### v1

# Probing cluster magnetism with embedded and background radio sources in Planck clusters

11institutetext: Dunlap Institute for Astronomy & Astrophysics, University of Toronto, 50 St. George Street, Toronto, ON M5S 3H4, Canada
11email: erik.osinga@utoronto.ca
22institutetext: Leiden Observatory, Leiden University, PO Box 9513, NL-2300 RA Leiden, The Netherlands 33institutetext: Minnesota Institute for Astrophysics, University of Minnesota, 116 Church St SE, Minneapolis, MN 55455, USA 44institutetext: Center for Astrophysics — Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 55institutetext: Department of Liberal Arts and Sciences, Berklee College of Music, 7 Haviland Street, Boston, MA 02215, USA 66institutetext: DIFA - Università di Bologna, via Gobetti 93/2, I-40129 Bologna, Italy 77institutetext: INAF - IRA, Via Gobetti 101, I-40129 Bologna, Italy; IRA - INAF, via P. Gobetti 101, I-40129 Bologna, Italy 88institutetext: Naval Research Laboratory, 4555 Overlook Avenue SW, Code 7213, Washington, DC 20375, USA 99institutetext: Institute for Astronomy, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK
Probing cluster magnetism with embedded and background radio sources in Planck clusters
E. Osinga 1122    R. J. van Weeren 22    L. Rudnick 33    F. Andrade-Santos 4455    A. Bonafede 6677    T. Clarke 88    K. Duncan 99    S. Giacintucci 88    H. J. A. Röttgering 22
(Received 2024-08-13; accepted Y)
Magnetic fields remain an enigmatic part of the content of galaxy clusters. Faraday rotation and depolarisation of extragalactic radio sources are useful probes, but the limited availability of polarised radio sources necessitates the stacking of clusters to study average magnetic field profiles and correlation scales. We recently presented a Karl G. Jansky Very Large Array survey of the 124 most massive Planck clusters at low redshift ($z<0.35$), where a clear depolarisation trend with the cluster impact parameter was found, with sources at smaller projected distances to the cluster centre showing more depolarisation. In this study, we combine the depolarisation information with the observed rotation measure (RM) and present an investigation into the average magnetic field properties of the sample, using both background sources and sources embedded in clusters. We observe a significant increase in the RM scatter, $\sigma_{\mathrm{RRM}}$, closer to the cluster centres. Averaging all 124 clusters, we find a scatter within $R_{\mathrm{500}}$ of $\sigma_{\mathrm{RRM}}=209\pm 37$ rad m-2, with background sources and cluster members showing similar values ($200\pm 33$ and $219\pm 66$ rad m-2, respectively). In the simple assumption of a uniform magnetic field with a single fluctuation scale $\Lambda_{c}$, this translates to an average magnetic field strength of $2\,(\Lambda_{c}/10\mathrm{kpc})^{-0.5}\,\mu$G. The profile of $\sigma_{\mathrm{RRM}}$ as a function of projected radius is inconsistent with a model that has a simple scaling $B\propto n_{e}^{\eta}$, with an observed deficit near the centre of clusters possibly caused by the fact that the highest RM sources near the centre of clusters are depolarised. Combining depolarisation and RM in a full forward model, we find that the magnetic field power spectrum roughly agrees with the Kolmogorov value, but that none of the Gaussian random field models can fully explain the observed relatively flat profiles. This implies that more sophisticated models of cluster magnetic fields in a cosmological context are needed.
Key Words.:
magnetic fields – polarization – galaxies: clusters: general – galaxies: clusters: intracluster medium – radiation mechanisms: non-thermal – methods: observational
1 Introduction
Galaxy clusters, the largest gravitationally bound structures in the Universe, harbour a rich variety of physical phenomena. Radio observations have revealed that clusters often show diffuse synchrotron emission that can span Mpc-sized regions, such as ‘radio halos’ (e.g. Bonafede et al., 2022) or ‘mega-halos’ (Cuciti et al., 2022), implying that clusters are filled with ultra-relativistic electrons and magnetic fields. The influence of the magnetic fields extends to particle acceleration models, radio synchrotron age estimates, the dynamics of the intracluster medium (ICM) and the transport of cosmic rays. Understanding the properties and origins of magnetic fields in clusters thus has broad importance (see Carilli & Taylor, 2002; Govoni & Feretti, 2004; Donnert et al., 2018, for reviews on magnetic fields in galaxy clusters).
The most promising tool to study magnetic fields is radio polarisation observations of Faraday rotation and depolarisation. A magnetised plasma such as the ICM causes a wavelength-dependent rotation of the polarisation angle. In general, the Faraday depth of a source is defined as (Burn, 1966; Brentjens & de Bruyn, 2005)
[…]
where $n_{e}$ is the electron density in parts per cm-3, B is the magnetic field in $\mu$Gauss and dr the infinitesimal path length increment along the line of sight (LOS) in kpc, and we define $\phi(\textbf{r})>0$ for the magnetic field pointing towards the observer. In the simple case of just one radio-emitting source along the line of sight, the Faraday depth is equal to the rotation measure (RM). With a combination of radio observations of RMs and X-ray observations of electron densities, it is thus possible to study the magnetic field properties of galaxy clusters.
Such studies are best done at low redshifts to maximise the cluster angular size because polarised radio sources are relatively rare (e.g. Rudnick & Owen, 2014). The most detailed analyses have been of the Coma Cluster (Bonafede et al., 2010) and Abell 2345 (Stuardi et al., 2021), where seven polarised radio sources were detected per cluster. The Coma Cluster magnetic field was found to agree with a Kolmogorov power spectrum with a central strength of 5 $\mu$G and a scaling of magnetic field energy density linearly proportional to the thermal gas density ($B^{2}\propto n_{e}$). The central magnetic field strength in Abell 2345 was found to be similar to the Coma Cluster, but with a magnetic field energy density that scales super-linearly instead (i.e. $B^{2}\propto n_{e}^{2}$). Several other low-redshift clusters have been analysed in polarisation (Murgia et al., 2004; Govoni et al., 2006; Guidetti et al., 2008; Govoni et al., 2010; Vacca et al., 2012; Govoni et al., 2017), with typically less than five polarised radio galaxies per study, resulting in large uncertainties on the magnetic field estimates (see e.g. Johnson et al., 2020, for a detailed discussion).
Because cluster magnetic fields are thought to be generally turbulent and disordered, the observed Faraday rotation is the result of a random walk process and thus a random variable. Accurate magnetic field estimates, therefore, require a statistical analysis probing many independent sight lines. Another potential problem is that polarised radio galaxies are often embedded in the cluster, and the degree to which the observed RM variations are caused by local interaction of the lobes with the ICM is debated (Rudnick & Blundell, 2003; Laing et al., 2008; Guidetti et al., 2012; Osinga et al., 2022). Such problems can be overcome by stacking clusters to increase the number of polarised radio sources located behind clusters and thus the number of independent sight lines through a cluster (Clarke et al., 2001; Bonafede et al., 2011; Böhringer et al., 2016; Stasyszyn & de los Rios, 2019; Osinga et al., 2022). Although stacking experiments have limited ability to probe differences between clusters, they are useful for obtaining average cluster magnetic field properties and are currently the only way to study cluster magnetic fields beyond the few nearest clusters.
We recently published the largest homogeneous stacking experiment using Karl G. Jansky Very Large Array (VLA) observations of 124 galaxy clusters (Osinga et al., 2022) selected from the Planck 2nd Sunyaev-Zeldovich Source Catalog (PSZ2) (Planck Collaboration et al., 2016). This study presented the first clear depolarisation trend tracing the radial profile of cluster magnetic fields using over 600 polarised radio sources. While depolarisation traces the smaller scale structure of the magnetic fields (i.e. few kpc), the larger scale structure can be inferred from the Faraday rotation of the radio sources. In this paper, we add the information from the Faraday rotation of the same sample of sources to study the large-scale properties of the magnetic fields in galaxy clusters. By jointly fitting both depolarisation and Faraday rotation, we aim to constrain the average magnetic field strength, scaling with density, and power spectrum. Cosmological calculations are performed assuming a flat $\Lambda$CDM model with $H_{0}=70$ kms-1Mpc-1, $\Omega_{m}=0.3$ and $\Omega_{\Lambda}=0.7$ and the $R_{\mathrm{500}}$ radius is the radius enclosing an overdensity of $500$ at the cluster redshift.
2 Chandra-Planck ESZ sample
The sample of galaxy clusters is a subset of 124 out of 165 clusters from the Chandra-Planck Legacy Program for Massive Clusters of Galaxies [^1] http://hea-www.cfa.harvard.edu/CHANDRA_PLANCK_CLUSTERS/ (Andrade-Santos et al., 2021) that have VLA observations presented by Osinga et al. (2022). The full details on the data reduction, polarised source identification and association, and determination of the polarisation properties are presented in the aforementioned paper, but we briefly summarise the important points here and highlight some improvements to the catalogue. All clusters are at low redshift ($z=0-0.35$) to maximise their apparent size. The analysis of the Chandra data and resulting electron density profiles is presented in Andrade-Santos et al. (2017).
Each of the 124 clusters was observed for $\sim 40$ min in the VLA L-band (1–2 GHz), resulting in typically 20–30 $\mu\mathrm{Jy\,beam}^{-1}$ noise levels at a resolution of 6–7′′  after data reduction. Polarised sources were identified using RM-synthesis (Brentjens & de Bruyn, 2005) and matched to total intensity components and optical counterparts. Optical counterparts were used to determine cluster membership, with spectroscopic redshifts where available and photometric redshifts otherwise. In total 6,807 and 819 source components were detected in total and polarised intensity respectively. Throughout this paper we will simply refer to these components as ‘sources’. We have fit the following model to the polarised intensity as a function of wavelength $\lambda$, which accounts for rotation and depolarisation (see Sokoloff et al., 1998, for details),
$P(\lambda^{2})=p_{0}I\exp(-2\Sigma^{2}_{\mathrm{RM}}\lambda^{4})\exp[2i(\chi_{0}+\phi\lambda^{2})],$ (2)
where $p_{0}$ denotes the intrinsic polarisation and $\Sigma^{2}_{\mathrm{RM}}$ the inferred variance of the RM distribution within the observing beam which models the depolarisation as a function of wavelength [^2] We note that Osinga et al. (2022) defined this as $\sigma^{2}_{\mathrm{RM}}$, but we use a different nomenclature here to avoid confusion with the scatter in RM between sources.. $\chi_{0}$ is the intrinsic polarisation angle, and $I$ denotes the total intensity model, which was assumed to be a simple power-law of the form $I(\nu)=I_{0}\nu^{\alpha}$ where $\nu$ is the frequency and $\alpha$ the spectral index.
We have improved the Monte Carlo Markov chain (MCMC) fitter used by Osinga et al. (2022) to now properly take into account the circular nature of $\chi_{0}$ during the fitting. Osinga et al. (2022) used a uniform prior on $\chi_{0}$, $\sim\mathcal{U}(0,\pi)$, which would cause the sampler in some cases to become trapped around the boundary values. We removed this prior on $\chi_{0}$ and fold the chain back into the range $[0,\pi)$ after the sampling is completed. We also calculate the mean and spread using circular statistics where we take into account the fact that the angles are distributed on the half-circle $\in$ $[0,\pi)$. In this way, the mean denotes the angle of the average vector on the unit circle, and the standard deviation is the spread in angles around the average vector. This agrees with the definition of the simple arithmetic mean and standard deviation when the angles are distributed away from the edges of the domain.
Table 1: Number counts of polarised sources and unique clusters in which they were detected for different subsets of the catalogue in Appendix C. Only clusters with X-ray data were forward-modelled (Sec. 4.4 & 4.5).
Subset Number of polarised sources Number of clusters
All 819 124
Used in analysis 610 117
Sources inside clusters 231 67
Sources behind clusters 363 101
$r/R_{\mathrm{500}}<1$ 261 68
$r/R_{\mathrm{500}}<1$ with X-ray 233 56
This paper is accompanied by an updated table of polarised components, shown in Appendix C. We note that this update mainly corrects the quoted mean and uncertainty of the intrinsic polarisation angle $\chi_{0}$ and most sources have similar best-fit RM and depolarisation parameters. This thus does not significantly impact the results. Finally, we used the same criteria for identifying bad fits as Osinga et al. (2022). All sources with a best-fit $\chi^{2}$ value that is $>5\sigma$ away from the theoretical distribution, and sources with low signal-to-noise polarised emission resulting in artificially large values of $\Sigma_{\mathrm{RM}}^{2}$ were flagged. This led to 195 bad fits among the 819 polarised sources identified in 124 galaxy clusters. After excluding three clusters (G121.11+57.01, G115.71+17.52, G139.59+24.18) due to a lack of accurate redshift estimates, a total of 610 sources across 117 clusters were available for subsequent analysis. Table 1 provides an overview of the number counts for various sub-samples and the unique clusters in which these sources were found, aiding in understanding the selection function of the defined subsets in the following analysis.
3 Methods
Figure 1: Simulated Faraday rotation (left) and depolarisation (right) screen for a random cluster from our sample as it would appear at 1.5 GHz with perfect sampling. The parameters are $B_{0}=5.0\,\mu$G, $n=2$, $\eta=0.5$ and the dashed and solid circles show $0.5\hbox{$R_{\mathrm{500}}$}$ and $1.0\hbox{$R_{\mathrm{500}}$}$, respectively.
3.1 Determining RM scatter
We first provide a brief note on nomenclature to ensure clarity regarding the terminology associated with various types of scatter. We adopt the following definitions throughout this paper:
  • $\delta$RM for the measurement uncertainty on RM.
  • $\sigma_{\mathrm{RM}}$ for the measurable scatter (i.e. standard deviation) in RM between sources.
  • $\Sigma_{\mathrm{RM}}$ for the inferred RM fluctuation across sources associated with beam depolarisation.
  • $\delta\sigma_{\mathrm{RM}}$ for uncertainty on the scatter in RM.
  • $\delta\Sigma_{\mathrm{RM}}$ for uncertainty on $\Sigma_{\mathrm{RM}}$.
[…]
$\begin{split}\mathrm{RM}_{\mathrm{obs^{\prime}}}&=\frac{\mathrm{RM}_{\mathrm{src}}}{(1+z_{\mathrm{intr}})^{2}}+\mathrm{RM}_{\mathrm{IGM}}+\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}\\ &\quad+\mathrm{RM}_{\mathrm{gal}}+\mathrm{RM}_{\mathrm{ion}}\end{split}$ (3)
The scatter in RM between sources, or RM variance, analogously is the sum of the variance of these individual terms, as they are independent contributions. The Galactic foreground can be largely removed by correcting the RMs for the Galactic contribution, which has been mapped most recently by Hutschenreuter et al. (2022). The measurement noise is known from the estimated uncertainties in RM from the MCMC spectral fitting. The most difficult components to isolate are the variance caused by the IGM and the intrinsic RM variance of radio sources. Studies using low RM variance sources (i.e. away from clusters and groups) that attempt to isolate the extragalactic from the Galactic and intrinsic scatter find that the extragalactic RM (rms) scatter is between $\sim$ 1 and 10 rad m-2 (Oppermann et al., 2015; Vernstrom et al., 2019; O’Sullivan et al., 2020; Pomakov et al., 2022).
In this study, we aim to separate the cluster RM scatter from the other sources of scatter. The cluster RM is often the largest term, as near the centre of clusters, RMs can be on the order of $10^{3}$ rad m-2. Meanwhile, the Galactic contribution to the RM is expected to be small since the cluster sample is selected from the PSZ2 survey and thus avoids the Galactic plane by design (Galactic latitude $|b|>14\hbox{${}^{\circ}$}$). We subtracted the Galactic contribution (on the order of $10^{1}$ rad m-2) using the recent map from Hutschenreuter et al. (2022), propagating the uncertainties in quadrature.
After subtraction of the Galactic contribution, we can define the residual RM as $\mathrm{RRM}=\mathrm{RM}_{\mathrm{obs^{\prime}}}-\mathrm{RM}_{\mathrm{gal}}$, and are thus left with
$\mathrm{RRM}=\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}+\delta\mathrm{RRM}+\mathrm{RM}_{\mathrm{extr}}$ (4)
where we defined the last term as the extrinsic RM, which captures the RM contribution that is not from the cluster or the measurement uncertainties. The extrinsic RM consists of components from the IGM, the radio source itself, and an ionospheric component, which are difficult to separate. However, we can estimate the scatter caused by $\mathrm{RM}_{\mathrm{extr}}$ through sources located far away from clusters. At large radii ($r>3R_{\mathrm{500}}$), the scatter is expected to be unrelated to the ICM, and caused mainly by the extrinsic scatter between sources and noise scatter due to measurement uncertainties. We find that the standard deviation of the observed RMs corrected for the Galactic contribution is equal to 13 $\pm$ 2 rad m-2 (see Fig. 18), with the measurement uncertainties accounting for 9 rad m-2. This gives an estimate of $\sigma_{\mathrm{RM,extr}}=9$ rad m-2, which is of the same order as the observed dispersion of extragalactic RMs (e.g. Schnitzeler, 2010; Vernstrom et al., 2019).
Using this estimate of the extrinsic scatter, we can define our estimator for the RM scatter induced by the ICM, in the observer frame. We define the RM scatter corrected for measurement errors and extrinsic scatter analogously to Dolag et al. (2001), as
${\sigma}_{\mathrm{RRM,cor^{\prime}}}=\sqrt{{\sigma}_{\mathrm{RRM}}^{2}-\frac{\sum_{i}^{N}(\delta\mathrm{RRM})_{i}^{2}}{N-1}-\sigma_{\mathrm{RM,extr}}^{2}},$ (5)
where $\sigma_{\mathrm{RRM}}$ is any estimator of the standard deviation, and the sum is taken over all sources $i=1$ to $i=N$ to account for varying measurement errors, $\delta\mathrm{RRM}$, per source, and $\sigma_{\mathrm{RM,extr}}=9$ rad m-2.
Finally, the cluster rest-frame RM scatter is higher than the observed RM scatter because we artificially measure smaller RMs (by a factor $(1+z)^{-2}$) due to cosmological redshift. This effect is difficult to account for exactly in a stacking experiment but can be approximated by using the median redshift of the cluster sample, weighted by the number of polarised sources per cluster. Finally, we thus define the corrected RRM scatter (standard deviation) as
$\sigma_{\mathrm{RRM,cor}}=\sigma_{\mathrm{RM,cor^{\prime}}}(1+\mathrm{med}[z_{\mathrm{cluster}}])^{2}.$ (6)
This estimator will be calculated in radial bins to measure the RM scatter profile.
3.2 Comparison to models
In the simple scenario of random magnetic field orientations in cells of size $\Lambda_{c}$ kpc, which have some magnetic field strength and electron density, the observed RM is the result of a random walk process. Because of the central limit theorem, the distribution of RMs is then expected to be a Gaussian distribution with zero mean, and variance given by (e.g. Murgia et al., 2004)
[…]
where dl is the infinitesimal path length increment along the line of sight (LOS) in kpc, $n_{e}$ is measured in cm-3 and $B_{\mathrm{||}}$ is the magnetic field strength parallel to the line of sight in $\mu$G.
In reality, the magnetic field structure will more closely resemble a random field with fluctuations on many spatial scales, and both the magnetic field strength and electron density will scale with radius. Thus comparing observations to more realistic scenarios requires simulated magnetic fields. We followed the approach explained in Section 4 of Osinga et al. (2022) to generate mock rotation measure and depolarisation images for all clusters in our sample which have X-ray observations available (99/124). An example of a mock RM and depolarisation image is shown in Figure 1 for a random cluster in our sample. In these models, the magnetic field is assumed to be a three-dimensional Gaussian random field with a single-power law spectrum characterised by the following parameters: $B_{0}$, $\eta$, $n$, $\Lambda_{\mathrm{min}}$ and $\Lambda_{\mathrm{max}}$. The first two denote the variables that parameterise the magnetic field, assumed to follow (e.g. Bonafede et al., 2010)
$B(r)=B_{0}\left(\frac{n_{\mathrm{e}}(r)}{\left<n_{\mathrm{e}}(0)\right>}\right)^{\eta},$ (8)
[…]
$|B_{k}|^{2}\propto k^{-n},$ (9)
between minimum and maximum fluctuations scales that are denoted by $\Lambda_{\mathrm{min}}$ and $\Lambda_{\mathrm{max}}$ in image space, respectively. In the picture of Kolmogorov turbulence, $n=11/3$, but this was found observationally to take values between $n=1$ and $n=4$ (Murgia et al., 2004; Govoni et al., 2006; Guidetti et al., 2008; Bonafede et al., 2010; Vacca et al., 2010, 2012; Govoni et al., 2017; Stuardi et al., 2021; Osinga et al., 2022). We have computed all models on $1024^{3}$ pixel grids, to simulate all clusters homogenously. Here, one pixel represents 3 kpc, and clusters are thus simulated out to about $1.5R_{\mathrm{500}}$, with a minimum fluctuation scale of $\Lambda_{\mathrm{min}}=6$ kpc. These models will be compared to observations in various ways, as detailed in the next section.
4 Results
4.1 Average magnetic field strength
In Figure 2, we plot the RRM as a function of projected distance to the nearest cluster centre (denoted as $a$ throughout this work). A clear trend is visible, with the scatter in RRM decreasing with a larger distance to the cluster centre. We determined the corrected RRM scatter for all sources within $1R_{\mathrm{500}}$, using Eq. 6, and find that the RRM scatter of cluster members and background sources are similar [^3] All uncertainties on $\sigma_{\mathrm{RRM}}$ are calculated by 1 000 bootstraps., being $200\pm 33$ and $219\pm 66$ rad m-2, respectively, as shown in Table 2. In reality, for a given magnetic field strength, the RRM scatter of background sources should be $\sqrt{2}$ times that of cluster members, as cluster members are on average located at the mid-plane of the cluster. However, cluster members are found preferentially at smaller radii, where the scatter in RM is larger due to generally larger magnetic field strength and electron densities.
Table 2: Corrected standard deviation of RM in rad m-2, as defined in Eq. 6 for different subsets and projected radii. We refer to Fig. 6 for the corrected IQR as a function of radius.
$<0.5R_{\mathrm{500}}$ $0.5-1.0R_{\mathrm{500}}$ $<R_{\mathrm{500}}$ $>R_{\mathrm{500}}$
All $244\pm 47$ $127\pm 30$ $209\pm 37$ $28\pm 5$
[…]
444
The mean value of the RMs agrees with zero as a function of radius, as shown in Figure 3, consistent with random magnetic field orientations along the line of sight. Assuming the simple random walk scenario denoted by Equation 7, we find that the most rudimentary estimate of the line-of-sight magnetic field strength is given by (e.g. Böhringer et al., 2016)
$\left(\frac{B_{\mathrm{||}}}{\mu\,G}\right)=2.46\left(\frac{\sigma_{\mathrm{RM}}/200}{\,\mathrm{rad\,m}^{-2}}\right)\left(\frac{n_{e}/10^{-3}}{\,\mathrm{cm}^{-3}}\right)\\ \left(\frac{\Lambda_{c}/10}{\,\mathrm{kpc}}\right)^{-1/2}\left(\frac{L/1000}{\,\mathrm{kpc}}\right)^{-1/2}$ (10)
where $L$ indicates the line-of-sight column length, which will be on average twice as large for background sources as cluster members. If we assume that cells are ordered on scales of $10$ kpc with an electron density of $10^{-3}\,\mathrm{cm}^{-3}$ (e.g. Böhringer et al., 2016), this reduces to
$B_{\mathrm{||}}=\frac{\sigma_{\mathrm{RM}}}{2.57\sqrt{L}}.$ (11)
If we assume approximately $L=1000$ kpc for cluster members and twice as large for background sources, we find magnetic field strengths averaged within $R_{\mathrm{500}}$ equal to 2-3 $\mu$G. Although the simple scenario suffices to give an order of magnitude estimate for the average magnetic field in galaxy clusters, it is clear from Figure 2 that the product of the magnetic field strength and electron density is not constant as a function of radius.
Figure 2: RRM as a function of normalised distance to the nearest cluster centre. The median uncertainty in RM is 11 rad m-2 for both cluster members and background sources. The plot is shown on a partially logarithmic scale in Appendix B to show low RM sources more clearly.
Figure 3: Mean of the RRM in bins of projected normalised distance to the nearest cluster centre. The horizontal error bars indicate the bin edges. Cluster members are shown in blue, background sources in orange, and the combined bins in black.
To sample a more physical property, we plot in Figure 4 the observed rotation measures as a function of ICM electron column density. This plot is less populated, as we now only show sources that are detected within a projected radius $a<2R_{\mathrm{500}}$, where the Chandra-derived column density values are reliable. The scatter in rotation measure significantly increases with increasing column density, and the preferential sampling of cluster members at high column densities (i.e. low radii) is clearly pronounced. Following Böhringer et al. (2016), we calculated the scatter in rotation measure in bins of column density, from which the average magnetic field strength along the line of sight can be calculated as
$\left(\frac{B_{\mathrm{||}}}{\mu\,G}\right)=3.801\times 10^{18}\left(\frac{\sigma_{\mathrm{RM}}}{\,\mathrm{rad\,m}^{-2}}\right)\left(\frac{N_{e}}{\,\mathrm{cm}^{-2}}\right)\left(\frac{L}{\Lambda_{c}}\right)^{1/2}.$ (12)
The RM scatter as a function of column density is shown in Figure 5. The bottom panel of Figure 5 shows the resulting magnetic field estimate from Equation 12. We find that the average magnetic field strength is $0.2\mu\mathrm{G}\sqrt{L/\Lambda_{c}}$, resulting in around $2\mu$G for typical values of $L=1000$ kpc and $\Lambda_{c}=10$ kpc (Böhringer et al., 2016).
Figure 4: RRM as a function of electron column density, for sources detected at $a<2R_{\mathrm{500}}$. The plot is shown on a partially logarithmic scale in Appendix B to show low RM sources more clearly.
Figure 5: Corrected standard deviation of RMs in bins of electron column density, for sources detected at $a<2R_{\mathrm{500}}$. The bins for the full sample have equal frequency with 71 sources per bin, while the bins for the sub-samples are defined in logarithmically spaced bins to sample similar densities. The horizontal error bars indicate the bins, and points are plotted at the median $N_{e}$. The bottom panel shows the magnetic field estimate assuming Equation 12, with a correlation length that is a factor 100 smaller than the line-of-sight distance, resulting in a mean strength of $\sim 2\leavevmode\nobreak\ \mu$G.
4.2 Radial profile
Figure 6: Corrected scatter of observed RRMs as a function of projected distance to the nearest cluster centre. The running scatter is calculated using a fixed sliding window of 50 points, the horizontal error bar indicates the median width of the sliding bin.
Though the average magnetic field strength is a useful property to constrain, cluster magnetic fields (and electron densities) show a radial decline. We calculated the scatter in rotation measure in a running window of 50 points as a function of radius. This was computed following Eq. 6, using the interquartile range (IQR) as a robust measure of ${\sigma}_{\mathrm{RRM}}$ (i.e. ${\sigma}_{\mathrm{RRM}}=\mathrm{IQR}/1.349$). A clear trend of the RM scatter being significantly higher closer to the cluster centre is present in Fig. 2, showing the effect of the turbulent magnetised ICM increasing in density, magnetic field strength or both  [^5] We present, in Appendix A an analysis into possible non-ICM effects or confounding variables that could cause such a scaling.. Background sources and cluster members show similar scatter profiles, indicating that they are largely tracing the same effect. However, background sources do show an increasing scatter above $a\sim 0.7R_{\mathrm{500}}$. We checked whether this can be attributed to a change in possible confounding variables in Appendix A (i.e. the uncertainty in RRM, galacic latitude of clusters, redshift of clusters, or polarisation fraction of radio sources), and found that this is the region where we start preferentially sampling high-mass clusters, in contrast to the inner region where low-mass clusters are preferentially sampled (see Fig. 13). As high-mass clusters provide larger column densities, and generally occupy denser parts of the cosmic large-scale structure with higher chances of intervening groups or dense filaments along the line-of-sight, it is likely that this is the cause of the rising RM scatter as a function of radius after $a\sim 0.7R_{\mathrm{500}}$.
[…]
In the simplified model of the turbulent ICM as laid out in Section 3, the variance in rotation measure as a function of projected distance should be proportional to the line-of-sight integral given in Equation 7. We can thus determine the magnitude of the 3D magnetic field fluctuations and the scaling between magnetic field and thermal electron density from the $\sigma_{\mathrm{RRM,cor}}$ profile. We sampled sources from our mock RM images created with $n=11/3$ (i.e. a Kolmogorov power spectrum), with the same radial sampling as the observed sources. We calculated the simulated scatter profile using the same sliding window, which is shown in Figure 7 for a magnetic field to electron density scaling of $\eta=0.5$ and $\eta=0.0$ (i.e. constant $B$). The canonical profile with $\eta=0.5$ is not a good fit to the data, as it underestimates the scatter near the outskirts and overestimates the scatter near the centre. The observed profile agrees more with a constant magnetic field, although given the smoothing scale of the sliding window, it is difficult to constrain the value of $\eta$ precisely. The best-fit central magnetic field strength likely lies between $1$ and $10$ $\mu G$, agreeing with the value from the depolarisation of radio sources (Osinga et al., 2022). We note that only using sources behind clusters gives similar results, as the scatter is similar as a function of radius. However, the relative flatness of the observed profile is caused by the cool-core cluster sample, as investigated in the next section.
Figure 7: Corrected scatter of RRMs as a function of projected distance to the nearest cluster centre for modelled clusters. The running scatter is calculated using a fixed sliding window of 50 points, with simulated profiles shown for $\eta=0.0$ and $\eta=0.5$.
4.3 Merging vs relaxed clusters
Differences between the magnetic field properties of relaxed and merging clusters have been tentatively observed in various studies (Bonafede et al., 2010; Stasyszyn & de los Rios, 2019; Osinga et al., 2022), but not yet clearly quantified. Following Osinga et al. (2022), we split our sample into merging and relaxed clusters based on the presence of a cool-core (CC) or absence of one (NCC). The corrected scatter in RM as a function of distance for the NCC and CC clusters is shown in Figure 8. There is a clear difference between the two samples, with the CC sample showing a relatively flat scatter profile, while the NCC sample shows significantly increasing scatter with lower projected radius. Similar behaviour was observed in the depolarisation of radio sources in the same sample, with CC clusters showing a flatter depolarisation profile (Osinga et al., 2022). Additionally, splitting the sample in background sources and cluster members (left panel and right panel) shows that both populations trace similar effects in NCC clusters whereas cluster sources show a significantly flatter profile in CC clusters. It is unlikely that this is an ICM effect, given that it would imply an increasing magnetic field as a function of radius. Given the anomalies in the CC sample, In Fig. 9, we compare the NCC clusters scatter profile to simulated scatter profiles. NCC clusters do show general agreement with a model with $\eta=0.5$ or $\eta=0.0$, although the drop in the central region $a<0.2R_{\mathrm{500}}$ cannot be explained by our model. This drop is observed in both the background population and sources inside clusters. It is possible that high RRM sources near the projected centre of clusters are depolarised, causing an observational bias against detecting high RRM sources near the centre. We exclude this $a<0.2R_{\mathrm{500}}$ region in the fit in Fig. 22, where we find that $\eta=0.0$ and $\eta=0.5$ have comparable $\chi^{2}$ values.
Figure 8: Corrected scatter of RRMs as a function of projected distance to the nearest cluster centre. Cool-core clusters are shown in blue, while non-cool-core clusters are shown in orange. The left panel shows background sources, while the right panel shows cluster members. The running scatter is calculated using a fixed sliding window of size 0.50, indicated by the horizontal errorbar.
Figure 9: Corrected scatter of RRMs observed in NCC clusters as a function of projected distance to the nearest cluster centre. The running scatter is calculated using a fixed sliding window of 50 points, the median width of the sliding window is indicated by the horizontal error bar. Best-fit simulated profiles are shown for $\eta=0.0$ and $\eta=0.5$.
4.4 Constraining the magnetic field power spectrum
Both the RM and depolarisation of sources scale similarly (although not identically) with the magnetic field strength, but differently with $n$ and $\Lambda_{\mathrm{max}}$, and therefore the ratio can be used to constrain $n$ and $\Lambda_{\mathrm{max}}$, relatively independent of the other parameters (e.g. Bonafede et al., 2010). We defined the ratio of RM to depolarisation as
$\frac{|\mathrm{RM}|}{1-\mathrm{DP}}.$ (13)
This ratio is also relatively independent of radial distance, as the depolarisation will increase (i.e. take lower values of DP) as the RM increases towards the cluster centre. We verified this with both Spearman and Pearson tests, which showed no significant correlation.
The median observed RM-DP ratio across all sources was found to be $113\pm 12$ rad m-2, with the uncertainty calculated by 1 00010001\,0001 000 bootstraps, with background sources and cluster members having similar values of $106\pm 17$ rad m-2 and $119\pm 14$ rad m-2, respectively. Figure 10 shows the median observed RM ratio compared with simulated values for sources sampled at similar positions in the simulated RM and depolarisation images. It is clear from the figure that models with most of the magnetic field energy on small scales (i.e. $n<3$) cannot reproduce the observed RM ratio, mainly because those models result in too low values of $|\mathrm{RM}|$ due to the rapidly fluctuating magnetic field along the line of sight. Instead, models with $n\geq 3$ provide a good fit for various values of the maximum correlation scale $\Lambda_{\mathrm{max}}$. Lowering $\Lambda_{\mathrm{max}}$ has an analogous effect to lowering $n$, namely decreasing the coherence length of the magnetic field along the line of sight. This thus results in a smaller average $|\mathrm{RM}|$ while the effect on depolarisation is less significant, as this is measured on scales below the observing beam (less than typically 15 kpc, although dependent on cluster redshift). Thus, the observed data best matches $n\geq 3$.
For a Kolmogorov spectrum ($n=3.67$) with typical values of the central magnetic field strength between $1-10\,\mu$G, the data is consistent with $\Lambda_{\mathrm{max}}>100$ kpc. Lower values of $\Lambda_{\mathrm{max}}$ would require significantly higher central magnetic field strengths.
Figure 10: RM ratio, as defined in Equation 13, as a function of magnetic field power spectrum index $n$. The median observed value is indicated by the black line, while the simulated values are indicated by the coloured points, with different markers for different magnetic field strengths. The uncertainty on the simulated values reflect the 16th and 84th percentile of 10 random initialisations.
4.5 RM and depolarisation forward model
Finally, we find the best model that reproduces the data by following the approach used in previous works such as Murgia et al. (2004); Bonafede et al. (2010); Govoni et al. (2017); Stuardi et al. (2021), directly comparing the simulated RM and depolarisation images to the observed data. For every source, we sample an equivalent source from the simulated clusters, and compute the expected RM and depolarisation. We minimize the difference between the simulated and observed radial scatter in RMs, and the depolarisation as a function of radius. We define the function to be minimized, $q$, as follows
$q=q_{\mathrm{depol}}+q_{\mathrm{RM}}=\\ \sum_{r}\left(\frac{\mathrm{DP}_{\mathrm{obs}}(r)-\left<\mathrm{DP}_{\mathrm{sim}}(r)\right>}{\mathrm{err}(\mathrm{DP}_{\mathrm{obs}})}\right)^{2}+\left(\frac{\sigma_{\mathrm{RRM,obs}}(r)-\left<\sigma_{\mathrm{RRM,sim}}(r)\right>}{\mathrm{err}(\sigma_{\mathrm{RRM,obs}})}\right)^{2},$ (14)
where the observables are calculated in running bins equivalent to Fig. 6. Simulated observables are averaged over 10 different random initialisations as denoted by the angle brackets. Because sources in this sample were found to be intrinsically depolarised with DP=0.92 at large radii (Osinga et al., 2022), we incorporated this into the simulated depolarisation. We fix the power spectrum to the Kolmogorov value of $n=3.67$ to reduce the computational burden, as Section 4.4 showed that the RM ratio is consistent only with models that have $n>3$.
Given the atypical behaviour of the CC sample, we forward model only the NCC sample, taking both behind and cluster members since the profiles are statistically similar (see Sec. 4.3). This smaller sample size consists of 182 sources detected at $a<1500$ kpc in 42 unique clusters. First, we investigate the best-fit models when fixing the magnetic field to electron density scaling to the typical value of $\eta=0.5$. Figures 11(a) and 11(b) show the values of $q_{\mathrm{depol}}$ and $q_{\mathrm{RM}}$ respectively, as a function of $B_{0}$ and $\Lambda_{\mathrm{max}}$. We find the best agreement for a model with $B_{0}=10\,\mu$G and a maximum correlation scale equal to $\Lambda_{\mathrm{max}}=100$ kpc for the depolarisation profile and $B_{0}=5\,\mu$G with $\Lambda_{\mathrm{max}}=1536$ kpc for the RRM scatter profile. The former is the best-fit model when combining the q values in a normalised form. Figure 12 shows the measured and simulated radial profiles for both best-fit models. Both models are unable to explain the data fully, with both the observed $\sigma_{\mathrm{RRM}}$ and the depolarisation profile being flatter than the simulated profiles.
Section 4.2 showed that the radial profile of $\sigma_{\mathrm{RRM}}$ preferred low values of $\eta\approx 0.0$. We thus also forward modelled with $\eta=0.0$, and show the q-values and best-fit forward models in Appendix B, Figures 23 and 24. Strictly by q-value, models with $\eta=0.0$ provide a better fit to the data than the $\eta=0.5$ models, but again none of the models provide a consistent fit to both the depolarisation and rotation measure scatter that we observed.
Figure 11: Values of $q_{\mathrm{depol}}$ (a) and $q_{\mathrm{RM}}$ (b) as defined in Equation 14 for combinations of $B_{0}$ and $\Lambda_{\mathrm{RM}}$. Models are simulated with a Kolmogorov power spectrum and $\eta=0.5$. The best-fit model is marked by a cross.
(a) Model that minimizes $q_{\mathrm{depol}}$.
(b) Model that minimizes $q_{\mathrm{RM}}$.
Figure 12: Comparison between observed and forward-modelled depolarisation and RM scatter for two models that minimize the q statistic as defined in Equation 14.
5 Discussion
Previous studies of samples of clusters have investigated either the statistical depolarisation (Bonafede et al., 2010; Osinga et al., 2022) or the rotation measure as a function of projected radius (Clarke et al., 2001; Böhringer et al., 2016; Stasyszyn & de los Rios, 2019). In this work, we have for the first time measured both the rotation measure and depolarisation in a consistent way for polarised radio sources along sight-lines of stacked galaxy clusters.
The previously largest statistical study of cluster magnetic fields was performed by Böhringer et al. (2016), who cross-matched 1773 clusters with the RM catalogue compiled by Kronberg & Newton-McGee (2011) to find 92 RM sight-lines within $R_{\mathrm{500}}$. We significantly increase the number statistics, detecting 261 radio sources with well-defined RM and depolarisation within $R_{\mathrm{500}}$ in only 124 clusters. Comparing Table 2 with Table 1 in Böhringer et al. (2016), we find significantly higher $\sigma_{\mathrm{RRM}}$ in all radial bins, except outside $R_{\mathrm{500}}$. The reason for this is likely twofold. First, there is a difference in the RM sensitivity as the Böhringer et al. (2016) is likely based on radio data with larger channel widths, preventing detection of high RMs (they find a maximum $|\mathrm{RM}|<700$ rad m-2). Second, the cluster samples are different, as Böhringer et al. (2016) quote a mean cluster mass of $3\times 10^{14}M_{\odot}$, while only 15 out of our 124 clusters have a mass below this value, and the mean mass of our sample is $5.7\times 10^{14}M_{\odot}$. Taking the different sampling into account by calculating the scatter as a function of electron column density in Section 4.1, we found an average magnetic field estimate of $3\,\mu$G in our sample of galaxy clusters, perfectly consistent with the findings of Böhringer et al. (2016). Outside the typical $R_{\mathrm{500}}$ radius of clusters, we observed a corrected standard deviation $\sigma_{\mathrm{RRM}}=28\pm 4$ rad m-2. This is lower than the value of $56\pm 8$ rad m-2 found by Böhringer et al. (2016) which did not correct for measurement uncertainties, but significantly higher than the expected intrinsic variation in the radio sources of $5-7$ rad m-2 (Schnitzeler, 2010). We thus observe an enhanced RM scatter beyond $R_{\mathrm{500}}$ as well.
Inside $R_{\mathrm{500}}$, the shape of the scatter profile was used to determine the magnetic field strength and electron density scaling, as shown in Figure 7. The RM scatter inside $R_{\mathrm{500}}$ showed a flat radial profile compared to expectations from a simple scaling of the magnetic field with the electron density of clusters. Plasma theories generally predict $\eta\geq 0.5$, with magnetic flux freezing giving $\eta=\frac{2}{3}$, adiabatic compression giving $\eta=1.0$ and dynamo models often predicting a constant magnetic energy density to thermal energy density ratio (i.e. $\eta=0.5$). Observationally, the best determined magnetic field profile is that of the Coma cluster, for which $\eta=0.4-0.7$ was found. Other studies using resolved (cluster) radio galaxies also find $\eta\geq 0.5$ (Murgia et al., 2004; Vacca et al., 2012; Govoni et al., 2017; Stuardi et al., 2021).
In contrast, we found that the best-fit $\sigma_{\mathrm{RRM}}(a)$ model preferred $\eta<0.5$, below the other experimental and theoretical values. However, statistical studies such as this one, that use RMs of unresolved radio sources to determine the RM scatter, suffer from a significant observational bias as sources with high RM values near the centre of clusters are likely to be depolarised and thus missed. The observed electron density profiles, shown in Fig 21, illustrate the problem as the electron densities rise strongly towards the core of clusters. Using Eq. 7, we expect $\sigma_{\mathrm{RRM}}$ to be  $\sim\,8$ times larger at $r=0.15R_{\mathrm{500}}$ than at $r=0.5R_{\mathrm{500}}$. Since the scatter observed at $0.5R_{\mathrm{500}}$ is on the order of $50$ rad m-2, we expect a scatter on the order of $400$ rad m-2 at $r=0.15R_{\mathrm{500}}$. Sources with such high RM scatter are quickly depolarised at the observed frequencies (Osinga et al., 2022), presenting a missing data problem. This artificially decreases the steepness of the $\sigma_{\mathrm{RRM}}$ profile, requiring low values of $\eta$ to match the data. In contrast, studies relying on resolved radio sources often probe $\sigma_{\mathrm{RRM}}$ on smaller scales (i.e. the size of the radio source) where fluctuations are expected to be significantly smaller and thus if a polarised radio source is detected, the scatter can be determined more accurately (e.g. Bonafede et al., 2010).
Other statistical studies probing many unresolved sightlines such as Böhringer et al. (2016); Stasyszyn & de los Rios (2019) have also not observed a strong increase of $\sigma_{\mathrm{RRM}}$ at low radii, implying the same observational bias. However, at higher frequencies, RM values as high as 10,000 rad m-2 have been observed in the centre of some clusters (e.g. Taylor & Perley, 1993). Higher frequency data that suffers less from depolarisation is thus needed to determine the value of $\eta$ accurately. This will be provided by, for example, the VLA Sky Survey (VLASS; Lacy et al., 2020) Depolarisation does not suffer as strongly from this observational bias at low radii, as upper limits on depolarisation fractions can still be set on sources that are significantly depolarised, given that they are detected at high signal-to-noise in total intensity. However, the depolarisation curves that we observed also do not agree with any of the Gaussian random field models explored in this paper, being generally flatter than models, as also seen in (Osinga et al., 2022).
Thus, we finally fit both the $\sigma_{\mathrm{RRM}}$ and depolarisation radial profile jointly by means of forward modelling the NCC clusters, since the CC clusters showed anomalous behaviour as explained below. None of the models could explain the RM and depolarisation data consistently, with observations finding a flatter profile than expected from the models. However, our models assume clusters are spherical over-densities without any surrounding large-scale structure, which is unrealistic since clusters will on average lie between one or two filaments (e.g. Pimbblet et al., 2004). The densities and possibly magnetic field strengths in the regions connecting the clusters to the filaments will be higher than our spherical $\beta$-like model, plausibly flattening the observed RM and depolarisation profiles. Comparisons with cosmological magneto-hydrodynamical (MHD) simulations are an obvious next step, where clusters are embedded in the large-scale structure around them (e.g. Marinacci et al., 2018).
Keeping these caveats in mind, we fixed the magnetic field to electron density scaling to the observationally best-determined value of $\eta=0.5$ (Bonafede et al., 2011) and found that the best-fit model had $B_{0}=5\,\mu$G and $\Lambda_{\mathrm{max}}\sim 100$ kpc. The central magnetic field strength is consistent with previous single object studies as well as statistical studies (Murgia et al., 2004; Govoni et al., 2006; Guidetti et al., 2008; Bonafede et al., 2010; Vacca et al., 2010, 2012; Govoni et al., 2017; Stuardi et al., 2021; Böhringer et al., 2016). However, the correlation scale is significantly larger than found in resolved cluster member studies that typically find values around $50$ kpc (Guidetti et al., 2008; Bonafede et al., 2010; Vacca et al., 2012; Govoni et al., 2017), although the resolved sources are often limited in size to less than a few hundred kiloparsec. Studies that use the brightness fluctuations or possibly polarised emission of radio halos to constrain the magnetic field power spectrum found results consistent with outer magnetic field fluctuation scales of $\sim 400$ kpc (Govoni et al., 2005; Vacca et al., 2010), agreeing with the 100 kpc estimate. Fluctuations on scales of more than a few hundred kpc are also expected theoretically, as the turbulent dynamo process thought to be responsible for magnetic field amplification in galaxy clusters is expected to occur on various scales, from less than a kpc up to a Mpc (Donnert et al., 2018).
In the forward modelling, we have decided to combine the information from cluster members and background sources to maximise the number statistics, as the $\sigma_{\mathrm{RRM}}$ profiles are statistically similar, and Osinga et al. (2022) showed that the depolarisation properties of cluster members and background sources are also similar. We checked for a biased contribution to $\sigma_{\mathrm{RRM}}$ from cluster members in Section 4.2 but found that, using the full sample, we could reject neither the null hypothesis that cluster members and background sources have similar scatter nor the null hypothesis that background sources show a scatter that is $\sqrt{2}$ times larger. This is in accordance with the interpretation from Osinga et al. (2022) that in large samples, cluster members and background sources trace similar ICM effects. However, it is also possible that we do not detect the $\sqrt{2}$ difference because cluster members have slightly enhanced RM scatter due to interaction with the ICM (e.g. Rudnick & Blundell, 2003). Because we could not reject either null hypothesis, a larger number of polarised radio sources is needed to observe whether background sources indeed show an RM scatter that is $\sqrt{2}$ times that of cluster members.
Finally, we checked for qualitative differences between merging and relaxed clusters. During cluster mergers, turbulence is injected on large scales, which can amplify magnetic fields, and drive large fluctuation scales (e.g. Vacca et al., 2018). Relaxed clusters, on the other hand, are expected to have smaller fluctuation scales, as the energy injected by a previous merger has dissipated through cascades to smaller and smaller scales. We found a significant difference between the merging and relaxed cluster samples. Merging clusters show higher values of $\sigma_{\mathrm{RRM}}$ in the region $0.5<r/R_{\mathrm{500}}$ than relaxed clusters, and sources inside relaxed clusters probe a flat $\sigma_{\mathrm{RRM}}$ profile as a function of radius. Higher RM scatter in merging clusters was also observed by Stasyszyn & de los Rios (2019), although with larger radial bin sizes. If a scaling between $B$ and $n_{e}$ is assumed, then CC clusters should show a steeper radial profile near the centre, as CC clusters often have higher central electron densities, as also shown in Fig 21. It is unclear what is causing the relative flatness of the $\sigma_{\mathrm{RRM}}$ profile in the sample of cluster members inside CC clusters, so CC clusters were not used for the full RM and depolarisation forward model.
6 Conclusion
This work has presented the continuation of the study presented by Osinga et al. (2022), where VLA 1-2 GHz polarisation observations of 124 massive Planck clusters were presented, and the depolarisation properties of polarised sources were investigated as a function of projected radius to the nearest cluster centre. We have incorporated the additional information from the best-fit RM and constrained cluster magnetic field properties by combining depolarisation and RM in a sample of clusters for the first time. We summarise the results of this work as follows:
  1. 1.
    We have clearly detected the increase of the scatter in rotation measure as a function of decreasing projected radius or increasing electron column density. Averaging all 124 clusters, we find a corrected scatter within $R_{\mathrm{500}}$ of $\sigma_{\mathrm{RM}}=209\pm 37$ rad m-2. The scatter outside of $R_{\mathrm{500}}$ was found to be $28\pm 5$ rad m-2, still significantly larger than our measurement uncertainties.
  2. 2.
    Assuming that magnetic fields fluctuate on a single characteristic length scale $\Lambda_{c}$ with a constant strength, the observed RM scatter agrees with an average magnetic field strength within $R_{\mathrm{500}}$ equal to $2\,(\Lambda_{c}/10\mathrm{kpc})^{-0.5}\,\mu$G.
  3. 3.
    Differences between the RM scatter of cluster members and background sources are expected from the differences in path lengths alone. However, using the full sample of sources, we could not reject the null hypothesis that cluster members sources show similar $\sigma_{\mathrm{RRM}}$ profiles as background sources, consistent with the result that background sources and cluster members also show similar depolarisation in the same sample of clusters (Osinga et al., 2022).
  4. 4.
    The profile of $\sigma_{\mathrm{RRM}}$ shows a significant decrease as a function of projected radius. However, the profile is flatter than expected from models that have a scaling of $B\propto n_{e}^{\eta}$ with $\eta=0.5$. We found that CC clusters show a flatter profile than NCC clusters, particularly when traced by cluster members.
  5. 5.
    Based on the ratio of RRM and depolarisation, models with magnetic field power spectral indices $n<3$ are strongly rejected by the data. This is consistent with expectations from Kolmogorov turbulence ($n=3.67$)
  6. 6.
    Jointly modelling both the depolarisation and rotation measure of sources in a forward modelling approach, we find that the observations cannot be explained by the simple model of a Gaussian random magnetic field that follows $B\propto n_{e}^{\eta}$. This implies that more advanced modelling of cluster magnetic fields is required.
In this work, we implicitly assumed that all clusters have the same magnetic field parameters, while in reality, this might be a function of dynamical state, mass, or redshift. The universality of cluster magnetic fields has not been thoroughly tested (e.g. Govoni et al., 2017) and might significantly affect our results. Future observations with more sensitive telescopes such as MeerKAT and the Square Kilometre Array could test this assumption by detecting enough polarised sightlines through single clusters such that stacking is not required.
Acknowledgements.
EO and RJvW acknowledge support from the VIDI research programme with project number 639.042.729, which is financed by the Netherlands Organisation for Scientific Research (NWO). EO thanks both Wout Goesaert and Joppe Swart for the careful looks at his code and Bryan Gaensler for providing comments on the manuscript. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. Basic research in radio astronomy at the U.S. Naval Research Laboratory is supported by 6.1 Base Funding. This paper has made use of observational material taken with an NRAO instrument. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. The Legacy Surveys project is honoured to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University This research has made use of the NASA/IPAC extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of NASA’s Astrophysics Data System (ADS). This research made use of Astropy, a community-developed core Python package for Astronomy. This research has made use of ChatGPT to restructure text and code.
Appendix A Confounding variables
The radial distribution of polarised sources has an important selection effect. The fact that clusters are observed with pointed observations of a finite primary beam size ($15^{\prime}$) means that there is a strong correlation between the projected radius of polarised sources and host cluster redshift, as shown in Fig. 13. This, combined with the fact that PSZ clusters, show a strong correlation between mass and redshift, with lower mass clusters only being found at low redshifts (e.g. Fig. 26 in Planck Collaboration et al. 2016), means that at small radii, the stacked sample is dominated by low-mass clusters, while at large radii the sample is dominated by high-mass clusters. Thus, even if clusters are self-similar in their thermal properties, at larger radii we will be probing preferentially larger column densities and thus expect larger RM scatter and depolarisation. Therefore, a simple analytical evaluation of Eq. 7 with a fixed path length does not follow the observed trend. However, if the radial profiles are compared to simulated profiles by forward modelling every cluster, integrating along appropriate path lengths, and sampling a simulated radio source at the same radial position, this is taken into account.
Figure 13: Distribution of all polarised radio sources in the (foreground) cluster vs projected radius plane. The black line shows the FWHM of the primary beam of the VLA L-band observations for a cluster of $R_{\mathrm{500}}=900$ kpc. The grey line shows the location where background sources start preferentially sampling higher redshift (higher mass) clusters, coinciding with the rising profile in Fig. 6.
We also investigate whether other variables than the projected radius correlate with increased scatter in RM. We found that the smallest background sources show the largest values of $|\mathrm{RRM}|$, as shown in Figure 14. Particularly, for the 29 sources with $\sqrt{\mathrm{Maj}\times\mathrm{Min}}<5$ arcseconds, the $|\mathrm{RRM}|$ rises rapidly. These sources show a median $|\mathrm{RRM}|=28$ rad m-2, while larger sources show a median of $|\mathrm{RRM}|=14.7$ rad m-2. This is not fully unexpected, as smaller sources probe a smaller (turbulent) foreground screen, and are thus less affected by smoothing. However, the results are not dominated by these 29 sources, as they are only a small subset of the whole sample, and the radial profile of the scatter is similar when excluding the smallest sources.
There is a strong correlation between the frational polarisation of radio sources and $|\mathrm{RRM}|$, as shown in Figure 15. This effect is particularly strong for cluster sources, which show significantly higher $|\mathrm{RRM}|$ at low polarisation fractions. Additionally, there are statistically more low polarization sources inside clusters than behind (2% KS probability of being the same). However, this effect is likely because the strongest depolarisation happens near the centre of the cluster (Osinga et al. 2022), where it is easier to detect cluster members than background sources. It is difficult to separate the effect of an intrinsic relation between low fractional polarisation of sources and higher rotation measures from the assumed effect of the ICM, but we can re-calculate the scatter profile using only sources with a fractional polarisation above 2%, where the relation between $|\mathrm{RRM}|$ and $p_{\mathrm{1.5GHz}}$ flattens. This is shown in Figure 16. The high fractional polarisation sources still show decreasing scatter as a function of radius, with similar profiles for cluster members and background sources. It is only the lowest fractional polarisation sources that show a significant difference between cluster members and background sources. Sources with low fractional polarisation are only found inside clusters when they are located near the centre. However, high fractional polarisation sources dominate the analysis, with 502 sources found with $p_{\mathrm{1.5GHz}}>0.02$ and only 107 sources found with $p_{\mathrm{1.5GHz}}<0.02$. The profile shown in Fig. 6 is thus dominated by high fractional polarisation sources, which explains that cluster and background sources follow similar profiles.
Finally, there is a strong correlation between $|\mathrm{RRM}|$ and Stokes I flux density for cluster members, where brighter sources in total intensity preferentially show higher RRM. Osinga et al. (2022) showed that we preferentially sample brighter sources near the cluster centre (i.e. there is a correlation between total flux density and projected radius from the cluster centre for cluster radio sources), which explains why the most centrally located, and thus brightest sources, correlate positively with $|\mathrm{RRM}|$. In conclusion, the aforementioned correlations are either not dominating the observed trend (i.e. in the case of small sources, or low fractional polarisation) or are likely a by-product of the observed ICM effect (i.e. in the case of the flux density correlation).
Figure 14: Running median of the absolute value of $\mathrm{RRM}$ as a function of source size. The distribution of sources in this parameter plane is plotted in the background and on the histograms beside the axes.
Figure 15: Running median of the absolute value of $\mathrm{RRM}$ as a function of polarisation fraction at 1.5 GHz. The distribution of sources in this parameter plane is plotted in the background and on the histograms beside the axes.
Figure 16: Corrected scatter in RRM as a function of projected radius, separately for sources with low ($p_{\mathrm{1.5GHz}}<0.02$, 107 sources) and high ($p_{\mathrm{1.5GHz}}>0.02$, 502 sources) fractional polarisation.
Figure 17: Absolute value of RRM as a function of Stokes I flux density.
Appendix B Alternative visualisations
In Fig. 18 we show a zoomed-in version of Fig. 2 beyond $a=3R_{\mathrm{500}}$, used to determine the scatter between sources far outside clusters. Additionally, to better visualise the low RM sources in our sample, we plot in Figures 19 and 20 the cluster-induced rotation measure on a logarithmic scale as a function of projected radius and column density, respectively. The electron density profiles of the clusters in our sample, as determined in Andrade-Santos et al. (2017), are shown in Figure 21.
Finally, the $q$ values of the full forward model for $\eta=0.0$ are shown in 23. The best-fit profiles for these fits are shown in Fig 24.
Figure 18: $\mathrm{RM}_{\mathrm{obs^{\prime}}}$ as a function of normalised distance to the nearest cluster centre for sources located far away from clusters.
Figure 19: RRM as a function of normalised distance to the nearest cluster centre. The y-axis is logarithmic for visualising low RM sources better than in Fig. 6. Uncertainties that are consistent with zero RM are plotted as downward-facing arrows.
Figure 20: RRM as a function of electron column density. The y-axis is logarithmic for visualising low RM sources better than in Fig. 2. Uncertainties that are consistent with zero RM are plotted as downward-facing arrows.
Figure 21: Average electron density profile as a function of normalised 3D radius for the 92 clusters that have Chandra X-ray observations. The mean profile is shown in the black line.
Figure 22: Like Fig. 9 but fit using only the data at $r>0.2R_{\mathrm{500}}$.
Figure 23: Values of $q_{\mathrm{depol}}$ (a) and $q_{\mathrm{RM}}$ (b) as defined in Equation 14 for combinations of $B_{0}$ and $\Lambda_{\mathrm{RM}}$. Models are simulated with a Kolmogorov power spectrum and $\eta=0.0$. The best-fit model is marked by a cross.
(a) Model that minimizes $q_{\mathrm{depol}}$.
(b) Model that minimizes $q_{\mathrm{RM}}$.
Figure 24: Comparison between observed and forward-modelled depolarisation and RM scatter for the model with $\eta=0.0$ that minimizes the q statistics as defined in Equation 14.
Appendix C Updated polarised source catalogue
Table 3: First 30 rows of the catalogue of 819 polarised radio sources that were detected in this work. The full table is available in electronic form at TODO
RA[deg] DEC[deg] Maj[′′] Min[′′] PA[deg] $p_{0}$ $\chi_{0}$[rad] RM[rad m-2] $\sigma_{\mathrm{RM}}$[rad m-2] $I_{0}$[mJy] $\alpha$ $\chi^{2}_{\mathrm{QU}}$ $z_{\mathrm{best}}$ $z_{\mathrm{best}}$ source $\theta_{p}$[arcmin] $r/R_{500}$ Cluster RAopt[deg] DECopt[deg] Multi component Flagged Note RRM[rad m-2]
$19.55453\pm 0.00009$ $-27.16767\pm 0.00010$ $10.4\pm 1.1$ $6.11\pm 0.35$ $136\pm 7$ $0.13^{+0.04}_{-0.03}$ $3.0^{+0.3}_{-0.3}$ $20^{+6}_{-5}$ $8^{+5}_{-3}$ $5.0^{+0.2}_{-0.2}$ $-0.5^{+0.1}_{-0.1}$ 61 $0.948\pm 0.077$ 2 $0.19$ $2.9$ G212.97-84.04 19.5606 -27.1622 False False $7^{+7}_{-7}$
$15.794198\pm 0.000035$ $-21.996243\pm 0.000030$ $7.03\pm 0.31$ $6.29\pm 0.24$ $117\pm 16$ $0.08^{+0.03}_{-0.02}$ $1.0^{+0.3}_{-0.3}$ $30^{+7}_{-6}$ $13^{+3}_{-3}$ $12.7^{+0.2}_{-0.2}$ $-0.88^{+0.04}_{-0.04}$ 52 - - $0.16$ $0.6$ G149.55-84.16 - - False False No optical counterpart. $24^{+8}_{-7}$
$41.4094860\pm 0.0000022$ $-20.6108492\pm 0.0000024$ $9.548\pm 0.021$ $8.841\pm 0.018$ $59.7\pm 1.2$ $0.104^{+0.001}_{-0.001}$ $0.465^{+0.008}_{-0.008}$ $13.6^{+0.1}_{-0.1}$ $2.4^{+0.4}_{-0.4}$ $149^{+2}_{-2}$ $-0.68^{+0.03}_{-0.03}$ 434 - - $0.21$ $3.5$ G205.07-62.94 - - False True No optical counterpart. $8^{+3}_{-3}$
$41.411690\pm 0.000009$ $-20.580840\pm 0.000007$ $9.18\pm 0.08$ $8.14\pm 0.06$ $101.3\pm 2.9$ $0.119^{+0.003}_{-0.003}$ $1.64^{+0.03}_{-0.03}$ $16.5^{+0.6}_{-0.6}$ $1.6^{+1.0}_{-0.6}$ $28.6^{+0.4}_{-0.4}$ $-0.83^{+0.04}_{-0.04}$ 144 - - $0.20$ $3.3$ G205.07-62.94 - - False True No optical counterpart. $10^{+3}_{-3}$
$40.48671\pm 0.00010$ $-28.63936\pm 0.00021$ $30.8\pm 1.9$ $10.4\pm 0.4$ $113.9\pm 3.3$ $0.28^{+0.02}_{-0.02}$ $0.37^{+0.07}_{-0.07}$ $8^{+1}_{-1}$ $7.2^{+1.0}_{-0.9}$ $16.1^{+0.5}_{-0.5}$ $-1.61^{+0.07}_{-0.08}$ 89 $0.238\pm 0.000$ 1 $0.11$ $1.6$ G222.97-65.69 40.4839 -28.6379 True True Radio Relic. $3^{+3}_{-3}$
$40.47708\pm 0.00020$ $-28.63432\pm 0.00015$ $20.7\pm 2.0$ $10.0\pm 0.6$ $124\pm 6$ $0.30^{+0.06}_{-0.05}$ $0.2^{+0.2}_{-0.2}$ $7^{+3}_{-3}$ $10^{+2}_{-2}$ $9.7^{+0.4}_{-0.4}$ $-2.4^{+0.1}_{-0.1}$ 80 $0.238\pm 0.000$ 1 $0.10$ $1.5$ G222.97-65.69 40.4839 -28.6379 True True Radio Relic. $2^{+4}_{-4}$
$40.48939\pm 0.00025$ $-28.65367\pm 0.00018$ $20.5\pm 2.3$ $14.8\pm 1.4$ $114\pm 16$ $0.14^{+0.01}_{-0.01}$ $1.9^{+0.2}_{-0.2}$ $6^{+3}_{-3}$ $2.2^{+1.8}_{-0.8}$ $9.6^{+0.4}_{-0.4}$ $-1.1^{+0.1}_{-0.1}$ 88 $0.238\pm 0.000$ 1 $0.11$ $1.6$ G222.97-65.69 40.4839 -28.6379 True True Radio Relic. $1^{+4}_{-4}$
$40.4898\pm 0.0005$ $-28.64934\pm 0.00008$ $46\pm 4$ $10.2\pm 0.5$ $96\pm 4$ $0.29^{+0.03}_{-0.03}$ $1.7^{+0.1}_{-0.1}$ $9^{+2}_{-2}$ $6^{+1}_{-1}$ $13.9^{+0.6}_{-0.6}$ $-1.9^{+0.1}_{-0.1}$ 131 $0.238\pm 0.000$ 1 $0.11$ $1.6$ G222.97-65.69 40.4839 -28.6379 True True Radio Relic. $5^{+3}_{-4}$
$40.47123\pm 0.00030$ $-28.63765\pm 0.00029$ $29.7\pm 3.2$ $17.2\pm 1.5$ $134\pm 11$ $0.19^{+0.04}_{-0.04}$ $1.6^{+0.2}_{-0.2}$ $7^{+3}_{-3}$ $9^{+2}_{-2}$ $20.3^{+0.7}_{-0.7}$ $-2.5^{+0.1}_{-0.1}$ 61 $0.238\pm 0.000$ 1 $0.09$ $1.4$ G222.97-65.69 40.4839 -28.6379 True True Radio Relic. $2^{+4}_{-5}$
$40.33575\pm 0.00011$ $-28.40285\pm 0.00006$ $10.4\pm 1.0$ $7.1\pm 0.4$ $104\pm 9$- -- -- -- -- -- - $1.135\pm 0.185$ 2 $0.25$ $3.5$ G222.97-65.69 - - True True No optical counterpart. --
$40.33588\pm 0.00004$ $-28.41099\pm 0.00004$ $10.4\pm 0.4$ $7.16\pm 0.18$ $138\pm 4$ $0.034^{+0.005}_{-0.005}$ $1.7^{+0.2}_{-0.2}$ $17^{+3}_{-3}$ $8^{+2}_{-1}$ $52.3^{+0.7}_{-0.7}$ $-0.96^{+0.03}_{-0.03}$ 85 $1.135\pm 0.185$ 2 $0.24$ $3.4$ G222.97-65.69 40.3352 -28.4072 True False $13^{+4}_{-4}$
$10.94092\pm 0.00004$ $-20.49858\pm 0.00004$ $9.5\pm 0.4$ $7.56\pm 0.25$ $143\pm 8$ $0.040^{+0.009}_{-0.006}$ $0.6^{+0.1}_{-0.1}$ $-10^{+3}_{-3}$ $7^{+3}_{-2}$ $17.1^{+0.3}_{-0.3}$ $-0.71^{+0.04}_{-0.04}$ 72 $0.414\pm 0.068$ 2 $0.14$ $2.3$ G106.73-83.22 10.9412 -20.4988 False False $-13^{+5}_{-5}$
$10.79686\pm 0.00014$ $-20.53360\pm 0.00005$ $22.8\pm 1.2$ $11.9\pm 0.4$ $84\pm 4$ $0.21^{+0.02}_{-0.02}$ $2.5^{+0.1}_{-0.1}$ $-6^{+2}_{-2}$ $7^{+1}_{-1}$ $15.0^{+0.5}_{-0.5}$ $-1.85^{+0.10}_{-0.10}$ 89 $0.281\pm 0.013$ 2 $0.09$ $1.4$ G106.73-83.22 10.7947 -20.5365 False False $-9^{+5}_{-5}$
$10.75281\pm 0.00012$ $-20.58744\pm 0.00006$ $18.1\pm 1.2$ $6.67\pm 0.17$ $155.7\pm 2.4$ $0.078^{+0.009}_{-0.008}$ $2.4^{+0.1}_{-0.1}$ $-11^{+2}_{-2}$ $6^{+2}_{-1}$ $18.3^{+0.4}_{-0.4}$ $-1.12^{+0.05}_{-0.05}$ 75 $0.729\pm 0.067$ 2 $0.09$ $1.4$ G106.73-83.22 10.7440 -20.5799 True False $-14^{+5}_{-5}$
$10.74349\pm 0.00005$ $-20.62051\pm 0.00012$ $14.2\pm 1.1$ $6.54\pm 0.22$ $159.5\pm 3.4$ $0.0014^{+0.0003}_{-0.0002}$ $0.2^{+0.2}_{-0.2}$ $17^{+4}_{-4}$ $8^{+3}_{-2}$ $479^{+5}_{-5}$ $-0.83^{+0.03}_{-0.02}$ 184 $0.281\pm 0.272$ 2 $0.10$ $1.5$ G106.73-83.22 10.7432 -20.6204 False True $14^{+6}_{-6}$
$10.743987\pm 0.000028$ $-20.57991\pm 0.00004$ $8.81\pm 0.35$ $7.17\pm 0.22$ $160\pm 7$ $0.0036^{+0.0008}_{-0.0006}$ $2.0^{+0.2}_{-0.2}$ $-157^{+4}_{-4}$ $6^{+4}_{-2}$ $134^{+2}_{-2}$ $-0.53^{+0.03}_{-0.03}$ 57 $0.729\pm 0.067$ 2 $0.10$ $1.6$ G106.73-83.22 10.7440 -20.5799 True False $-160^{+6}_{-6}$
$10.73185\pm 0.00011$ $-20.57074\pm 0.00017$ $14.4\pm 1.4$ $11.5\pm 1.0$ $172\pm 19$ $0.09^{+0.02}_{-0.01}$ $1.1^{+0.2}_{-0.2}$ $-3^{+4}_{-3}$ $6^{+3}_{-2}$ $9.5^{+0.3}_{-0.3}$ $-0.60^{+0.08}_{-0.08}$ 73 $0.729\pm 0.067$ 2 $0.11$ $1.8$ G106.73-83.22 10.7440 -20.5799 True False $-6^{+5}_{-5}$
10.8157 -20.5492 8.2 8.2 0 $0.19^{+0.05}_{-0.04}$ $3.0^{+0.2}_{-0.2}$ $-19^{+4}_{-4}$ $6^{+3}_{-2}$ $3.2^{+0.3}_{-0.3}$ $-2.2^{+0.2}_{-0.2}$ 71 $0.603\pm 0.041$ 2 $0.07$ $1.0$ G106.73-83.22 10.8167 -20.5483 False False $-22^{+6}_{-6}$
$46.89273\pm 0.00006$ $-28.725980\pm 0.000029$ $7.1\pm 0.5$ $5.03\pm 0.22$ $106\pm 7$ $0.4^{+0.3}_{-0.3}$ $0^{+1}_{-1}$ $83^{+74}_{-75}$ $41^{+5}_{-8}$ $11.8^{+0.2}_{-0.2}$ $-0.57^{+0.04}_{-0.04}$ 120 $0.270\pm 0.020$ 2 $0.13$ $1.9$ G223.91-60.09 46.8925 -28.7259 False False $76^{+74}_{-75}$
$46.83815\pm 0.00017$ $-28.68291\pm 0.00004$ $13.7\pm 1.5$ $6.18\pm 0.30$ $95\pm 5$ $0.03^{+0.02}_{-0.01}$ $2.5^{+0.3}_{-0.3}$ $40^{+9}_{-7}$ $12^{+6}_{-3}$ $20.1^{+0.3}_{-0.3}$ $-0.85^{+0.04}_{-0.04}$ 78 $0.383\pm 0.035$ 2 $0.07$ $1.1$ G223.91-60.09 46.8365 -28.6812 True False $33^{+10}_{-8}$
$46.83576\pm 0.00015$ $-28.67959\pm 0.00021$ $17.2\pm 2.2$ $6.00\pm 0.30$ $145\pm 5$ $0.07^{+0.05}_{-0.02}$ $2.5^{+0.5}_{-0.5}$ $43^{+19}_{-16}$ $24^{+7}_{-4}$ $25.1^{+0.4}_{-0.3}$ $-0.76^{+0.03}_{-0.03}$ 71 $0.383\pm 0.035$ 2 $0.06$ $1.1$ G223.91-60.09 46.8365 -28.6812 True False $35^{+20}_{-16}$
$46.83814\pm 0.00005$ $-28.68534\pm 0.00005$ $7.2\pm 0.5$ $5.38\pm 0.26$ $138\pm 9$ $0.04^{+0.51}_{-0.02}$ $2.1^{+0.6}_{-0.6}$ $79^{+18}_{-48}$ $17^{+2283}_{-5}$ $11.9^{+0.2}_{-0.2}$ $-0.71^{+0.04}_{-0.04}$ 54 $0.383\pm 0.035$ 2 $0.07$ $1.1$ G223.91-60.09 46.8365 -28.6812 True True $72^{+19}_{-48}$
$357.900782\pm 0.000032$ $-25.938221\pm 0.000018$ $7.68\pm 0.28$ $5.76\pm 0.15$ $102\pm 5$ $0.038^{+0.010}_{-0.007}$ $0.4^{+0.2}_{-0.2}$ $-26^{+4}_{-4}$ $7^{+3}_{-2}$ $16.5^{+0.2}_{-0.2}$ $-0.55^{+0.04}_{-0.04}$ 63 $0.187\pm 0.133$ 3 $0.14$ $2.0$ G034.03-76.59 357.9006 -25.9388 False False $-34^{+6}_{-6}$
$156.14934\pm 0.00006$ $-27.20391\pm 0.00004$ $7.9\pm 0.6$ $6.10\pm 0.31$ $112\pm 10$ $0.04^{+0.02}_{-0.01}$ $1.7^{+0.4}_{-0.4}$ $48^{+12}_{-10}$ $16^{+5}_{-3}$ $13.3^{+0.2}_{-0.2}$ $-0.39^{+0.04}_{-0.04}$ 76 $0.258\pm 0.163$ 3 $0.17$ $2.4$ G266.84+25.07 156.1492 -27.2033 False False $49^{+21}_{-20}$
$156.02742\pm 0.00005$ $-27.331459\pm 0.000032$ $8.4\pm 0.4$ $7.05\pm 0.27$ $90\pm 10$ $0.06^{+0.01}_{-0.01}$ $0.4^{+0.2}_{-0.2}$ $-127^{+6}_{-6}$ $15^{+3}_{-2}$ $10.8^{+0.2}_{-0.2}$ $-0.78^{+0.04}_{-0.04}$ 95 - - $0.08$ $1.1$ G266.84+25.07 - - False False No optical counterpart. $-124^{+21}_{-21}$
$155.935295\pm 0.000027$ $-27.34246\pm 0.00004$ $8.82\pm 0.34$ $7.21\pm 0.22$ $11\pm 7$ $0.07^{+0.01}_{-0.01}$ $1.4^{+0.2}_{-0.2}$ $-132^{+7}_{-7}$ $20^{+2}_{-2}$ $17.3^{+0.3}_{-0.3}$ $-0.55^{+0.03}_{-0.03}$ 113 - - $0.08$ $1.2$ G266.84+25.07 - - False False No optical counterpart. $-129^{+19}_{-19}$
$155.723060\pm 0.000033$ $-27.180613\pm 0.000032$ $9.01\pm 0.29$ $8.66\pm 0.26$ $116\pm 33$ $0.025^{+0.004}_{-0.003}$ $0.9^{+0.1}_{-0.1}$ $-9^{+3}_{-2}$ $5^{+2}_{-2}$ $43.3^{+0.6}_{-0.6}$ $-1.01^{+0.04}_{-0.04}$ 138 - - $0.24$ $3.4$ G266.84+25.07 - - False False No optical counterpart. $-11^{+25}_{-25}$
$172.93573\pm 0.00008$ $-19.87149\pm 0.00008$ $8.0\pm 0.7$ $7.7\pm 0.6$ $(1.4\pm 0.8)\times 10^{2}$ $0.3^{+0.6}_{-0.2}$ $0^{+1}_{-1}$ $-142^{+958}_{-739}$ $43^{+192880647}_{-17}$ $2.6^{+0.2}_{-0.2}$ $-0.7^{+0.1}_{-0.1}$ 80 $0.291\pm 0.025$ 3 $0.09$ $1.3$ G278.60+39.17 172.9325 -19.8793 True True $-132^{+958}_{-739}$
$172.932507\pm 0.000035$ $-19.87992\pm 0.00007$ $13.5\pm 0.6$ $8.47\pm 0.23$ $20\pm 4$ $0.044^{+0.005}_{-0.005}$ $2.7^{+0.2}_{-0.2}$ $-0^{+3}_{-3}$ $4^{+2}_{-1}$ $10.0^{+0.3}_{-0.3}$ $-0.95^{+0.06}_{-0.06}$ 152 $0.291\pm 0.025$ 3 $0.08$ $1.2$ G278.60+39.17 172.9325 -19.8793 True True $10^{+11}_{-11}$
$172.93171\pm 0.00005$ $-19.88390\pm 0.00005$ $9.1\pm 0.5$ $8.00\pm 0.35$ $139\pm 16$ $0.08^{+0.02}_{-0.02}$ $2.8^{+0.3}_{-0.3}$ $-115^{+7}_{-7}$ $14^{+3}_{-2}$ $4.8^{+0.2}_{-0.2}$ $-0.69^{+0.08}_{-0.08}$ 141 $0.291\pm 0.025$ 3 $0.08$ $1.2$ G278.60+39.17 172.9325 -19.8793 True False $-105^{+13}_{-13}$
 [^6] The columns are identical to Osinga et al. (2022), except the newly added column RRM, which is the observed RM corrected for the Galactic value. If uncertainties are not available, they are not given.

### v2

@@
 11email: erik.osinga@utoronto.ca
-22institutetext: Leiden Observatory, Leiden University, PO Box 9513, NL-2300 RA Leiden, The Netherlands 33institutetext: Minnesota Institute for Astrophysics, University of Minnesota, 116 Church St SE, Minneapolis, MN 55455, USA 44institutetext: Center for Astrophysics — Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 55institutetext: Department of Liberal Arts and Sciences, Berklee College of Music, 7 Haviland Street, Boston, MA 02215, USA 66institutetext: DIFA - Università di Bologna, via Gobetti 93/2, I-40129 Bologna, Italy 77institutetext: INAF - IRA, Via Gobetti 101, I-40129 Bologna, Italy; IRA - INAF, via P.
-Gobetti 101, I-40129 Bologna, Italy 88institutetext: Naval Research Laboratory, 4555 Overlook Avenue SW, Code 7213, Washington, DC 20375, USA 99institutetext: Institute for Astronomy, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK
+22institutetext: Leiden Observatory, Leiden University, PO Box 9513, NL-2300 RA Leiden, The Netherlands 33institutetext: Minnesota Institute for Astrophysics, University of Minnesota, 116 Church St SE, Minneapolis, MN 55455, USA 44institutetext: Center for Astrophysics — Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 55institutetext: Department of Liberal Arts and Sciences, Berklee College of Music, 7 Haviland Street, Boston, MA 02215, USA 66institutetext: Clay Center Observatory, Dexter Southfield, 20 Newton Street, Brookline, MA 02445, USA 77institutetext: DIFA - Università di Bologna, via Gobetti 93/2, I-40129 Bologna, Italy 88institutetext: INAF - IRA, Via Gobetti 101, I-40129 Bologna, Italy; IRA - INAF, via P.
+Gobetti 101, I-40129 Bologna, Italy 99institutetext: Naval Research Laboratory, 4555 Overlook Avenue SW, Code 7213, Washington, DC 20375, USA 1010institutetext: Institute for Astronomy, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK
 Probing cluster magnetism with embedded and background radio sources in Planck clusters
-E. Osinga 1122    R. J. van Weeren 22    L. Rudnick 33    F. Andrade-Santos 4455    A. Bonafede 6677    T. Clarke 88    K. Duncan 99    S. Giacintucci 88    H. J. A. Röttgering 22
+E. Osinga 1122    R. J. van Weeren 22    L. Rudnick 33    F. Andrade-Santos 445566    A. Bonafede 7788    T. Clarke 99    K. Duncan 1010    S. Giacintucci 99    H. J. A. Röttgering 22
 (Received 2024-08-13; accepted Y)
-Magnetic fields remain an enigmatic part of the content of galaxy clusters.
-Faraday rotation and depolarisation of extragalactic radio sources are useful probes, but the limited availability of polarised radio sources necessitates the stacking of clusters to study average magnetic field profiles and correlation scales.
+Magnetic fields remain an elusive part of the content of galaxy clusters.
+Faraday rotation and depolarisation of extragalactic radio sources are useful probes, but the limited availability of polarised radio sources necessitates the stacking of clusters to study average magnetic field properties.
 We recently presented a Karl G.
-Jansky Very Large Array survey of the 124 most massive Planck clusters at low redshift ($z<0.35$), where a clear depolarisation trend with the cluster impact parameter was found, with sources at smaller projected distances to the cluster centre showing more depolarisation.
+Jansky Very Large Array survey of the 124 most massive Planck clusters at low redshift ($z<0.35$), finding a clear depolarisation trend with the cluster impact parameter, with sources at smaller projected distances to the cluster centre showing more depolarisation.
 In this study, we combine the depolarisation information with the observed rotation measure (RM) and present an investigation into the average magnetic field properties of the sample, using both background sources and sources embedded in clusters.
@@
 Averaging all 124 clusters, we find a scatter within $R_{\mathrm{500}}$ of $\sigma_{\mathrm{RRM}}=209\pm 37$ rad m-2, with background sources and cluster members showing similar values ($200\pm 33$ and $219\pm 66$ rad m-2, respectively).
-In the simple assumption of a uniform magnetic field with a single fluctuation scale $\Lambda_{c}$, this translates to an average magnetic field strength of $2\,(\Lambda_{c}/10\mathrm{kpc})^{-0.5}\,\mu$G.
+In the simple assumption of a uniform amplitude magnetic field with a single fluctuation scale $\Lambda_{c}$, this translates to an average magnetic field strength of $2\,(\Lambda_{c}/10\mathrm{kpc})^{-0.5}\,\mu$G.
 The profile of $\sigma_{\mathrm{RRM}}$ as a function of projected radius is inconsistent with a model that has a simple scaling $B\propto n_{e}^{\eta}$, with an observed deficit near the centre of clusters possibly caused by the fact that the highest RM sources near the centre of clusters are depolarised.
@@
 Understanding the properties and origins of magnetic fields in clusters thus has broad importance (see Carilli & Taylor, 2002; Govoni & Feretti, 2004; Donnert et al., 2018, for reviews on magnetic fields in galaxy clusters).
-The most promising tool to study magnetic fields is radio polarisation observations of Faraday rotation and depolarisation.
-A magnetised plasma such as the ICM causes a wavelength-dependent rotation of the polarisation angle.
+The most promising tool to study magnetic fields is radio polarisation observations.
+A magnetised plasma such as the ICM causes a wavelength-dependent rotation of the polarisation angle (Faraday rotation) and depolarisation.
 In general, the Faraday depth of a source is defined as (Burn, 1966; Brentjens & de Bruyn, 2005)
@@
 In the simple case of just one radio-emitting source along the line of sight, the Faraday depth is equal to the rotation measure (RM).
-With a combination of radio observations of RMs and X-ray observations of electron densities, it is thus possible to study the magnetic field properties of galaxy clusters.
+With a combination of RMs from radio observations and electron densities from X-ray observations, it is thus possible to study the magnetic field properties of galaxy clusters.
 Such studies are best done at low redshifts to maximise the cluster angular size because polarised radio sources are relatively rare (e.g.
@@
 The Coma Cluster magnetic field was found to agree with a Kolmogorov power spectrum with a central strength of 5 $\mu$G and a scaling of magnetic field energy density linearly proportional to the thermal gas density ($B^{2}\propto n_{e}$).
-The central magnetic field strength in Abell 2345 was found to be similar to the Coma Cluster, but with a magnetic field energy density that scales super-linearly instead (i.e.
+The central magnetic field strength in Abell 2345 was found to be similar to the Coma Cluster, but with a magnetic field energy density that scales more steeply instead (i.e.
 $B^{2}\propto n_{e}^{2}$).
@@
 This study presented the first clear depolarisation trend tracing the radial profile of cluster magnetic fields using over 600 polarised radio sources.
-While depolarisation traces the smaller scale structure of the magnetic fields (i.e.
-few kpc), the larger scale structure can be inferred from the Faraday rotation of the radio sources.
+While depolarisation traces the smaller scale structure of the magnetic fields on the scale of the resolving beam (usually kpc), the larger scale structure (i.e.
+along the line of sight, usually Mpc) can be inferred from the Faraday rotation of the radio sources.
 In this paper, we add the information from the Faraday rotation of the same sample of sources to study the large-scale properties of the magnetic fields in galaxy clusters.
@@
 Optical counterparts were used to determine cluster membership, with spectroscopic redshifts where available and photometric redshifts otherwise.
+The full details on redshift determination are given in Osinga et al.
+(2022), where Eq.
+25-26 were used to divide sources into background and cluster members.
 In total 6,807 and 819 source components were detected in total and polarised intensity respectively.
 Throughout this paper we will simply refer to these components as ‘sources’.
-We have fit the following model to the polarised intensity as a function of wavelength $\lambda$, which accounts for rotation and depolarisation (see Sokoloff et al., 1998, for details),
+We have fit the following model to the polarised intensity as a function of wavelength $\lambda$, which accounts for rotation and depolarisation
 $P(\lambda^{2})=p_{0}I\exp(-2\Sigma^{2}_{\mathrm{RM}}\lambda^{4})\exp[2i(\chi_{0}+\phi\lambda^{2})],$ (2)
@@
 $\chi_{0}$ is the intrinsic polarisation angle, and $I$ denotes the total intensity model, which was assumed to be a simple power-law of the form $I(\nu)=I_{0}\nu^{\alpha}$ where $\nu$ is the frequency and $\alpha$ the spectral index.
+Equation 2 assumes that (see Sokoloff et al., 1998, for details):
+  1. 1.
+    The dominant mechanism causing Faraday rotation is an external Faraday screen (i.e. a non-emitting foreground magnetised plasma, such as the ICM).
+  2. 2.
+    The foreground screen magnetic field changes on scales (at least) smaller than the observing beam, which causes Faraday dispersion inside the beam, resulting in depolarisation.
+  3. 3.
+    The magnetic field fluctuates as a Gaussian random field, implying its scatter can be completely captured in the parameter $\Sigma_{\mathrm{RM}}$.
+It has been shown (Osinga et al., 2022) that this model accurately fits 75% of the polarised radio sources detected in this sample, and we discard sources that do not show a proper fit (see below).
 We have improved the Monte Carlo Markov chain (MCMC) fitter used by Osinga et al.
@@
 In this way, the mean denotes the angle of the average vector on the unit circle, and the standard deviation is the spread in angles around the average vector.
-This agrees with the definition of the simple arithmetic mean and standard deviation when the angles are distributed away from the edges of the domain.
+This agrees with the definition of the simple arithmetic mean and standard deviation when the angles are distributed away from the edges of the domain (see e.g.
+Mardia & Jupp, 2009, for more details on circular statistics).
 Table 1: Number counts of polarised sources and unique clusters in which they were detected for different subsets of the catalogue in Appendix C.
-Only clusters with X-ray data were forward-modelled (Sec.
-4.4 & 4.5).
 Subset Number of polarised sources Number of clusters
@@
 $r/R_{\mathrm{500}}<1$ with X-ray 233 56
+ [^3] Only clusters with X-ray data were forward-modelled (Sec. 4.4 & 4.5).
 This paper is accompanied by an updated table of polarised components, shown in Appendix C.
-We note that this update mainly corrects the quoted mean and uncertainty of the intrinsic polarisation angle $\chi_{0}$ and most sources have similar best-fit RM and depolarisation parameters.
+We note that this update mainly corrects the quoted mean and uncertainty of the intrinsic polarisation angle $\chi_{0}$ and most sources have similar best-fit RM ($\phi$) and depolarisation ($\Sigma_{\mathrm{RM}}$) parameters to within a few per cent.
 This thus does not significantly impact the results.
@@
   • $\Sigma_{\mathrm{RM}}$ for the inferred RM fluctuation across sources associated with beam depolarisation.
-  • $\delta\sigma_{\mathrm{RM}}$ for uncertainty on the scatter in RM.
+  • $\delta\sigma_{\mathrm{RM}}$ for uncertainty on the scatter in RM ($\sigma_{\mathrm{RM}}$).
   • $\delta\Sigma_{\mathrm{RM}}$ for uncertainty on $\Sigma_{\mathrm{RM}}$.
 […]
-$\begin{split}\mathrm{RM}_{\mathrm{obs^{\prime}}}&=\frac{\mathrm{RM}_{\mathrm{src}}}{(1+z_{\mathrm{intr}})^{2}}+\mathrm{RM}_{\mathrm{IGM}}+\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}\\ &\quad+\mathrm{RM}_{\mathrm{gal}}+\mathrm{RM}_{\mathrm{ion}}\end{split}$ (3)
+$\begin{split}\mathrm{RM}_{\mathrm{obs^{\prime}}}&=\frac{\mathrm{RM}_{\mathrm{src}}}{(1+z_{\mathrm{intr}})^{2}}+\mathrm{RM}_{\mathrm{IGM}}+\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}\\ &\quad+\mathrm{RM}_{\mathrm{gal}}+\mathrm{RM}_{\mathrm{ion}},\end{split}$ (3)
+where $\mathrm{RM}_{\mathrm{IGM}}$ is the integrated contribution of the IGM, which itself has a redshift dependence (c.f.
+Amaral et al., 2021, Eq.
+2) that is however not relevant for this work as we regard it as a single nuisance parameter.
 The scatter in RM between sources, or RM variance, analogously is the sum of the variance of these individual terms, as they are independent contributions.
+The measurement noise is known from the estimated uncertainties in RM from the MCMC spectral fitting (we use the $16$th and $84$th percentile of the posterior, see Osinga et al., 2022, for details).
 The Galactic foreground can be largely removed by correcting the RMs for the Galactic contribution, which has been mapped most recently by Hutschenreuter et al.
 (2022).
-The measurement noise is known from the estimated uncertainties in RM from the MCMC spectral fitting.
 The most difficult components to isolate are the variance caused by the IGM and the intrinsic RM variance of radio sources.
@@
 After subtraction of the Galactic contribution, we can define the residual RM as $\mathrm{RRM}=\mathrm{RM}_{\mathrm{obs^{\prime}}}-\mathrm{RM}_{\mathrm{gal}}$, and are thus left with
-$\mathrm{RRM}=\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}+\delta\mathrm{RRM}+\mathrm{RM}_{\mathrm{extr}}$ (4)
-where we defined the last term as the extrinsic RM, which captures the RM contribution that is not from the cluster or the measurement uncertainties.
+$\mathrm{RRM}=\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}+\mathrm{RM}_{\mathrm{extr}}$ (4)
+where we defined the last term as the extrinsic RM, which captures the RM contribution that is not from the cluster or the measurement uncertainties ($\delta\mathrm{RRM}$).
 The extrinsic RM consists of components from the IGM, the radio source itself, and an ionospheric component, which are difficult to separate.
@@
 3.2 Comparison to models
-In the simple scenario of random magnetic field orientations in cells of size $\Lambda_{c}$ kpc, which have some magnetic field strength and electron density, the observed RM is the result of a random walk process.
+In the simple scenario of random magnetic field directions in cells of size $\Lambda_{c}$ kpc, which have some uniform magnetic field strength and uniform electron density, the observed RM is the result of a random walk process.
 Because of the central limit theorem, the distribution of RMs is then expected to be a Gaussian distribution with zero mean, and variance given by (e.g.
@@
 (2022) to generate mock rotation measure and depolarisation images for all clusters in our sample which have X-ray observations available (99/124).
+We assumed that the clusters have spherical and smooth electron density profiles, with best-fits to the de-projected electron density as a function of radius taken from Andrade-Santos et al.
+(2017), which are shown in Fig 21.
+In reality, there are likely density fluctuations in the ICM.
+However, these are difficult to measure and on the order of only a few per cent (e.g.
+Schuecker et al., 2004; Churazov et al., 2012; Sanders & Fabian, 2012).
+Magnetic field fluctuations are expected to be stronger.
+Therefore, we have modelled the magnetic field as a Gaussian random field.
 An example of a mock RM and depolarisation image is shown in Figure 1 for a random cluster in our sample.
+To compare the mock images to observations, we sample a mock source for every observed source at the same projected radius.
+If the source was classified as a background source, the integration is done over the full grid, while for a cluster member, the integration is done assuming the source lies at the midplane of the cluster.
 In these models, the magnetic field is assumed to be a three-dimensional Gaussian random field with a single-power law spectrum characterised by the following parameters: $B_{0}$, $\eta$, $n$, $\Lambda_{\mathrm{min}}$ and $\Lambda_{\mathrm{max}}$.
-The first two denote the variables that parameterise the magnetic field, assumed to follow (e.g.
-Bonafede et al., 2010)
+The first two denote the variables that parameterise the magnetic field strength, assumed to follow
 $B(r)=B_{0}\left(\frac{n_{\mathrm{e}}(r)}{\left<n_{\mathrm{e}}(0)\right>}\right)^{\eta},$ (8)
 […]
+The last three parameters encode the power spectrum of the magnetic field:
 $|B_{k}|^{2}\propto k^{-n},$ (9)
-between minimum and maximum fluctuations scales that are denoted by $\Lambda_{\mathrm{min}}$ and $\Lambda_{\mathrm{max}}$ in image space, respectively.
-In the picture of Kolmogorov turbulence, $n=11/3$, but this was found observationally to take values between $n=1$ and $n=4$ (Murgia et al., 2004; Govoni et al., 2006; Guidetti et al., 2008; Bonafede et al., 2010; Vacca et al., 2010, 2012; Govoni et al., 2017; Stuardi et al., 2021; Osinga et al., 2022).
-We have computed all models on $1024^{3}$ pixel grids, to simulate all clusters homogenously.
-Here, one pixel represents 3 kpc, and clusters are thus simulated out to about $1.5R_{\mathrm{500}}$, with a minimum fluctuation scale of $\Lambda_{\mathrm{min}}=6$ kpc.
-These models will be compared to observations in various ways, as detailed in the next section.
+[…]
 4 Results
@@
 6, and find that the RRM scatter of cluster members and background sources are similar [^5] All uncertainties on $\sigma_{\mathrm{RRM}}$ are calculated by 1 000 bootstraps., being $200\pm 33$ and $219\pm 66$ rad m-2, respectively, as shown in Table 2.
-In reality, for a given magnetic field strength, the RRM scatter of background sources should be $\sqrt{2}$ times that of cluster members, as cluster members are on average located at the mid-plane of the cluster.
-However, cluster members are found preferentially at smaller radii, where the scatter in RM is larger due to generally larger magnetic field strength and electron densities.
-Table 2: Corrected standard deviation of RM in rad m-2, as defined in Eq. 6 for different subsets and projected radii. We refer to Fig. 6 for the corrected IQR as a function of radius.
+Assuming that clusters are spherically symmetrical on average, or at least have no bias towards our direction, and that cluster members are on average located at the mid-plane of the cluster, we expect the RRM scatter of background sources to be at least $\sqrt{2}$ times that of cluster members at the same projected radius.
+However, cluster members are found preferentially at smaller radii (e.g.
+Osinga et al., 2022, Fig.
+B1), where the scatter in RM is larger due to generally larger magnetic field strength and electron densities (e.g.
+Bonafede et al., 2010).
+Table 2: Corrected standard deviation of RM in rad m-2, as defined in Eq. 6 for different subsets and projected radii.
 $<0.5R_{\mathrm{500}}$ $0.5-1.0R_{\mathrm{500}}$ $<R_{\mathrm{500}}$ $>R_{\mathrm{500}}$
@@
 […]
-444
+ [^6] We refer to Fig. 6 for the corrected IQR as a function of radius.
 The mean value of the RMs agrees with zero as a function of radius, as shown in Figure 3, consistent with random magnetic field orientations along the line of sight.
-Assuming the simple random walk scenario denoted by Equation 7, we find that the most rudimentary estimate of the line-of-sight magnetic field strength is given by (e.g.
-Böhringer et al., 2016)
-$\left(\frac{B_{\mathrm{||}}}{\mu\,G}\right)=2.46\left(\frac{\sigma_{\mathrm{RM}}/200}{\,\mathrm{rad\,m}^{-2}}\right)\left(\frac{n_{e}/10^{-3}}{\,\mathrm{cm}^{-3}}\right)\\ \left(\frac{\Lambda_{c}/10}{\,\mathrm{kpc}}\right)^{-1/2}\left(\frac{L/1000}{\,\mathrm{kpc}}\right)^{-1/2}$ (10)
+Assuming the simple random walk scenario denoted by Equation 7, we find that the most rudimentary estimate of the line-of-sight magnetic field strength is given by
+$\left(\frac{B_{\mathrm{||}}}{\mu\,G}\right)=2.46\left(\frac{\sigma_{\mathrm{RM}}/200}{\,\mathrm{rad\,m}^{-2}}\right)\left(\frac{n_{e}/10^{-3}}{\,\mathrm{cm}^{-3}}\right)^{-1}\\ \left(\frac{\Lambda_{c}/10}{\,\mathrm{kpc}}\right)^{-1/2}\left(\frac{L/1000}{\,\mathrm{kpc}}\right)^{-1/2}$ (10)
 where $L$ indicates the line-of-sight column length, which will be on average twice as large for background sources as cluster members.
@@
 Cluster members are shown in blue, background sources in orange, and the combined bins in black.
-To sample a more physical property, we plot in Figure 4 the observed rotation measures as a function of ICM electron column density.
+As we have information on the radial shape of the electron densities, we plot in Figure 4 the observed rotation measures as a function of ICM electron column density.
 This plot is less populated, as we now only show sources that are detected within a projected radius $a<2R_{\mathrm{500}}$, where the Chandra-derived column density values are reliable.
@@
 We verified this with both Spearman and Pearson tests, which showed no significant correlation.
-The median observed RM-DP ratio across all sources was found to be $113\pm 12$ rad m-2, with the uncertainty calculated by 1 00010001\,0001 000 bootstraps, with background sources and cluster members having similar values of $106\pm 17$ rad m-2 and $119\pm 14$ rad m-2, respectively.
-Figure 10 shows the median observed RM ratio compared with simulated values for sources sampled at similar positions in the simulated RM and depolarisation images.
+The median observed ratio of ${|\mathrm{RM}|}/(1-\mathrm{DP})$ across all sources was found to be $113\pm 12$ rad m-2, with the uncertainty calculated by 1 00010001\,0001 000 bootstraps, with background sources and cluster members having similar values of $106\pm 17$ rad m-2 and $119\pm 14$ rad m-2, respectively.
+Figure 10 shows the median ratio compared with simulated values for sources sampled at similar positions in the simulated RM and depolarisation images.
 It is clear from the figure that models with most of the magnetic field energy on small scales (i.e.
@@
 (2016); Stasyszyn & de los Rios (2019) have also not observed a strong increase of $\sigma_{\mathrm{RRM}}$ at low radii, implying the same observational bias.
-However, at higher frequencies, RM values as high as 10,000 rad m-2 have been observed in the centre of some clusters (e.g.
-Taylor & Perley, 1993).
+However, at higher frequencies, RM values as high as 10,000 rad m-2 have been observed in the centre of some clusters [^8] Depending on the electron density and magnetic field structure in the cluster, Taylor & Perley (1993) determined that a magnetic field strength of 30 $\mu$G was required in case of the Hydra A cluster.
+(e.g.
+Taylor & Perley, 1993; Baidoo et al., 2023).
 Higher frequency data that suffers less from depolarisation is thus needed to determine the value of $\eta$ accurately.
@@
 The densities and possibly magnetic field strengths in the regions connecting the clusters to the filaments will be higher than our spherical $\beta$-like model, plausibly flattening the observed RM and depolarisation profiles.
-Comparisons with cosmological magneto-hydrodynamical (MHD) simulations are an obvious next step, where clusters are embedded in the large-scale structure around them (e.g.
+Additionally, our models do not account for any fluctuations in the electron density with a similar or different power spectrum than that of the magnetic field.
+Although there are indications from simulations that the magnetic field and thermal density power spectra could be similar at small scales (e.g.
+Domínguez-Fernández et al., 2019), this is difficult to constrain observationally and expected to be on the order of a few per cent (e.g.
+Churazov et al., 2012).
+Additional density fluctuations would generally increase the modelled RM scatter, and thus require slightly lower magnetic field strengths to fit observations.
+Finally, the assumption that the magnetic fields are Gaussian is in contrast with expectations if the fields are generated by the fluctuation dynamo mechanism, which predicts spatially intermittent fields with non-Gaussian statistics (e.g.
+Schekochihin et al., 2004; Vazza et al., 2018; Seta et al., 2020; Kopyev et al., 2022) Comparisons with cosmological magneto-hydrodynamical (MHD) simulations are an obvious next step, with realistic turbulent gas profiles, potential non-Gaussian magnetic fields and clusters that are embedded in the large-scale structure around them (e.g.
 Marinacci et al., 2018).
@@
 It is unclear what is causing the relative flatness of the $\sigma_{\mathrm{RRM}}$ profile in the sample of cluster members inside CC clusters, so CC clusters were not used for the full RM and depolarisation forward model.
+In the future, exploring a two-component magnetic field model for CC clusters might alleviate the inconsistency, analogous to two-component models that are used to fit CC cluster thermal electron density profiles (e.g.
+Vikhlinin et al., 2006).
 6 Conclusion
@@
     Jointly modelling both the depolarisation and rotation measure of sources in a forward modelling approach, we find that the observations cannot be explained by the simple model of a Gaussian random magnetic field that follows $B\propto n_{e}^{\eta}$.
-This implies that more advanced modelling of cluster magnetic fields is required.
+Although $\eta=0$ is slightly preferred over $\eta=0.5$, more advanced modelling of cluster magnetic fields is required.
 In this work, we implicitly assumed that all clusters have the same magnetic field parameters, while in reality, this might be a function of dynamical state, mass, or redshift.
@@
 Future observations with more sensitive telescopes such as MeerKAT and the Square Kilometre Array could test this assumption by detecting enough polarised sightlines through single clusters such that stacking is not required.
+Such observations should take into account that Equation 2 might not hold for all sources, as diffuse radio emission from clusters (e.g.
+van Weeren et al., 2019) might induce complexity into the Faraday spectra.
+Data availability
+Table 3 is available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
 Acknowledgements.
+We thank the referee, Amit Seta, for a thorough report that has improved the quality of this manuscript.
 EO and RJvW acknowledge support from the VIDI research programme with project number 639.042.729, which is financed by the Netherlands Organisation for Scientific Research (NWO).
+RJvW acknowledges support from the ERC Starting Grant ClusterWeb 804208.
 EO thanks both Wout Goesaert and Joppe Swart for the careful looks at his code and Bryan Gaensler for providing comments on the manuscript.
@@
 2 beyond $a=3R_{\mathrm{500}}$, used to determine the scatter between sources far outside clusters.
-Additionally, to better visualise the low RM sources in our sample, we plot in Figures 19 and 20 the cluster-induced rotation measure on a logarithmic scale as a function of projected radius and column density, respectively.
+We find no significant correlation between RRM and projected radius for sources far outside the central regions, although statistically significant peaks in RRM have been found at large radii in a larger sample of RRMs around cosmic overdensities (Anderson et al.
+2024).
+To better visualise the low RM sources in our sample, we plot in Figures 19 and 20 the cluster-induced rotation measure on a logarithmic scale as a function of projected radius and column density, respectively.
 The electron density profiles of the clusters in our sample, as determined in Andrade-Santos et al.
 (2017), are shown in Figure 21.
-Finally, the $q$ values of the full forward model for $\eta=0.0$ are shown in 23. The best-fit profiles for these fits are shown in Fig 24.
+Finally, the $q$ values of the full forward model for $\eta=0.0$ are shown in 23.
+The best-fit profiles for these fits are shown in Fig 24.
 Figure 18: $\mathrm{RM}_{\mathrm{obs^{\prime}}}$ as a function of normalised distance to the nearest cluster centre for sources located far away from clusters.
@@
 Appendix C Updated polarised source catalogue
-Table 3: First 30 rows of the catalogue of 819 polarised radio sources that were detected in this work. The full table is available in electronic form at TODO
+Table 3: First 30 rows of the catalogue of 819 polarised radio sources that were detected in this work.
 RA[deg] DEC[deg] Maj[′′] Min[′′] PA[deg] $p_{0}$ $\chi_{0}$[rad] RM[rad m-2] $\sigma_{\mathrm{RM}}$[rad m-2] $I_{0}$[mJy] $\alpha$ $\chi^{2}_{\mathrm{QU}}$ $z_{\mathrm{best}}$ $z_{\mathrm{best}}$ source $\theta_{p}$[arcmin] $r/R_{500}$ Cluster RAopt[deg] DECopt[deg] Multi component Flagged Note RRM[rad m-2]
@@
 If uncertainties are not available, they are not given.
+The full table is available in electronic form at the CDS.